### v1

# Probabilistic Approaches to The Energy Equality in Forced Surface Quasi-Geostrophic Equations

Abstract.
We explore probabilistic approaches to the deterministic energy equality for the forced Surface Quasi-Geostrophic (SQG) equation on a torus. First, we prove the zero-noise dynamical large deviations for a corresponding stochastic SQG equation, where the lower bound matches the upper bound on a certain closure of the weak-strong uniqueness class for the deterministic forced SQG equation. Furthermore, we show that the energy equality for the deterministic SQG equation holds on arbitrary time-reversible subsets of the domain, where we match the upper bound and the lower bound. Conversely, the violation of the deterministic energy equality breaks the lower bound of large deviations. These results extend the existing techniques in Gess, Heydecker, and the second author [GHW24] to generalized Sobolev spaces with negative indices. Finally, we provide an analysis of the restricted quasi-potential and prove a conditional equivalence compared to the rate function of large deviations for the Gaussian distribution. This suggests a potential connection between non-Gaussian large deviations in equilibrium for the stochastic SQG equation and the open problem regarding the uniqueness of the deterministic SQG equation.
Key words and phrases:
Surface Quasi-Geostrophic equation, large deviations, energy equality.
2010 Mathematics Subject Classification:
60H15; 60F10; 35Q35
1. Introduction
This paper is dedicated to exploring probabilistic approaches to the deterministic energy equality for a forced dissipative Surface Quasi-Geostrophic (SQG) equation on the torus $\mathbb{T}^{2}=\mathbb{R}^{2}/\mathbb{Z}^{2}$:
$\partial_{t}\theta=$ $-\Lambda^{2\alpha}\theta-u_{\theta}\cdot\nabla\theta+\Lambda^{2\beta}g,$ (1.1)
$u_{\theta}=$ $R^{\perp}\theta=(-R_{2}\theta,R_{1}\theta),$
where $\Lambda=(-\Delta)^{1/2},R_{j}$ $(j=1,2)$ is the $j$-th Riesz transform and $g\in L^{2}([0,T];L^{2}(\mathbb{T}^{2}))$. The parameters $\alpha$ and $\beta$ will be discussed within the following ranges:
$\alpha\in(0,1/2),\beta=\alpha/2,\ \ \text{and}\ \ \alpha\in[1/2,1),\beta=\alpha/2\text{ or }\alpha/2+1/4.$ (1.2)
A key element of these probabilistic approaches is the interpretation of (1.1) as the so-called skeleton equation of the following stochastic SQG equation:
$\partial_{t}\theta_{\varepsilon}=$ $-\Lambda^{2\alpha}\theta_{\varepsilon}-u_{\theta_{\varepsilon}}\cdot\nabla\theta_{\varepsilon}+\varepsilon^{1/2}\Lambda^{2\beta}\xi_{\delta(\varepsilon)},$ (1.3)
$u_{\theta_{\varepsilon}}=$ $R^{\perp}\theta_{\varepsilon}=(-R_{2}\theta_{\varepsilon},R_{1}\theta_{\varepsilon}),$
within the framework of large deviations theory. Here, $\xi_{\delta(\varepsilon)}$ is a spatial regularization of space-time white noise $\xi$. We will explore the connections between dynamical large deviations of (1.3) as $(\varepsilon,\delta(\varepsilon))\rightarrow(0,0)$ and the validity of the deterministic energy equality in (1.1). Additionally, we will highlight the potential relationships between large deviations in equilibrium for (1.3) and the uniqueness problem for (1.1). Furthermore, when $\varepsilon=0$ in (1.3), a scaling argument suggests that the fractional dissipation index $\alpha=1/2$ belongs to the critical regime, $\alpha>1/2$ falls within the subcritical regime and $\alpha\in(0,1/2)$ stands for the supercritical regime [Res95]. Hence the ranges in (1.2) cover subcritical, critical, and supercritical cases.
Regarding the deterministic PDE (1.1), the initial data is considered to be an $L^{2}(\mathbb{T}^{2})$-function with zero mean, and the Leray-Hopf solution theory (see [Res95, Theorem 3.1], analogous to the Navier-Stokes equations) is employed. One of the key features of the Leray solutions is the so-called energy inequality. However, it is unknown whether the equality case of the energy inequality
$\frac{1}{2}\|\theta(T)\|_{L^{2}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\|\Lambda^{\alpha}\theta\|_{L^{2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\|\theta(0)\|_{L^{2}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\langle\Lambda^{2\beta}\theta,g\rangle\,\mathrm{d}s$ (1.4)
holds for a Leray solution of (1.1). Analogous to the celebrated result of Lions-Ladyzhenskaya for three-dimensional Navier-Stokes equations, it is possible to verify the energy equality for the SQG equation under some regularity conditions, see for example [Dai17]. On the other hand, the uniqueness problem of (1.1) remains open as well, with various regularity conditions specified to show uniqueness. We refer readers to [Res95, Section 3.4], [DC06] and the references therein for more details. To the best of our knowledge, in the study of SQG equation, there are no direct relationships between the regularity conditions for the energy equality and those for uniqueness. This is one of the motivations for this work, and we point out that a direct relationship will be provided in this context based on the perspective of probability. More precisely, this is linked to the study of dynamical large deviations of (1.3). Furthermore, this unveils a sufficient condition for proving the energy equality (1.4), although it is unclear how to verify such a condition.
The stochastic PDE (1.3) (proposed by Totz [Tot20] and Hofmanová et al. [HLZZ24]) is an SQG equation driven by additive noise with small intensity $\varepsilon$ and small correlation $\delta(\varepsilon)$. The consideration of such noise follows from the principles of fluctuating hydrodynamics theory. Generally speaking, the SQG equation can be regarded as a toy model for understanding the regularity of the Navier-Stokes equations. We point out that an important fluctuating Navier-Stokes model known as the Landau-Lifshitz-Navier-Stokes (or Navier-Stokes-Fourier system for the full equations) is governed by the fluctuation-dissipation principle, a fundamental concept in the theory of fluctuating hydrodynamics (see [Ö05]). The stochastic SQG equation (1.3) could be considered as a stochastic toy model for the Landau-Lifshitz-Navier-Stokes equations. In particular, the choice of $\beta=\frac{\alpha}{2}$, corresponding to the fractional dissipation $-\Lambda^{2\alpha}$, exhibits a structural resemblance to the fluctuation-dissipation principle. Regarding (1.3), a dynamical large deviation principle is proved, in a joint scaling regime $(\varepsilon,\delta(\varepsilon))\rightarrow(0,0)$, with initial data allowing for fluctuations as well. Moreover, we provide an analysis on the identification of the quasi-potential
$\mathcal{U}(\phi)=\frac{1}{2}\inf\Big{\{}\int^{0}_{-\infty}\|\partial_{t}\tilde{\theta}+\Lambda^{2\alpha}\tilde{\theta}+R^{\perp}\tilde{\theta}\cdot\nabla\tilde{\theta}\|_{H^{-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s:\ -\tilde{\theta}(-\cdot)\in\mathcal{A}_{0},\ \tilde{\theta}(0)=\phi\Big{\}},$ (1.5)
where $\phi\in H^{\alpha-2\beta}(\mathbb{T}^{2})$ and $\mathcal{A}_{0}$ is a restricted domain related to the dynamical large deviations result. We will specify $\mathcal{A}_{0}$ later on.
Here, we emphasize that in the study of large deviations for (1.3), the initial data could also be considered in $\dot{H}^{-1/2}(\mathbb{T}^{2})$. In such a lower regularity regime, a commutator estimate associated with the Riesz transform is employed to make the nonlinear term well-defined. This distinguishes the technique used in proving large deviations for SQG equations from that for three-dimensional Landau-Lifshitz-Navier-Stokes equations established by Gess, Heydecker, and the second author [GHW24]. Further comments regarding the technique and the physical background for such an $\dot{H}^{-1/2}(\mathbb{T}^{2})$ framework will be provided later on.
A theoretical basis for considering the fluctuating hydrodynamics equation is that the Gibbs measure preserves the dynamics invariant informally [Spo91]. As long as we replace the correlated noise $\xi_{\delta(\varepsilon)}$ with a space-time white noise $\xi$, and choose $\beta=\frac{\alpha}{2}$, we are able to see that the $L^{2}(\mathbb{T}^{2})$-cylindrical Gaussian measure $\mathcal{G}(0,\varepsilon I/2)$ is informally invariant for the dynamics, where $I$ is the identity operator on $L^{2}(\mathbb{T}^{2})$, and the dynamics are time-reversible with respect to this measure. A rigorous mathematical approach can be found within the energy solution framework [Tot20], where the existence of stationary energy solutions with Gaussian distribution is demonstrated. This raises an important question: when regularizing space-time white noise and studying fluctuations in a more regular space under the scaling regimes $(\varepsilon,\delta(\varepsilon))$, can the Gaussian equilibrium feature be preserved in terms of the rate function for large deviations?
We highlight the work [BC17] and [CP22], where the authors established a large deviation principle for the invariant measure of the two-dimensional Navier-Stokes equation and showed that the rate function is governed by its quasi-potential, which coincides with the rate function for Gaussian large deviations. Inspired by their work, one can investigate large deviations in equilibrium of regularized stochastic PDEs (1.3). Whether an asymptotic Gaussian or non-Gaussian equilibrium feature arises depends on whether the corresponding quasi-potential coincides with the Gaussian rate function. This motivates us to investigate the identification of the quasi-potential (1.5) for the SQG equation.
Roughly speaking, the main results in this paper provide the relationships described by the following picture,
with
(UL): Weak-strong uniqueness implies a restricted dynamical large deviation principle;
(LE): The energy equality holds on any time-reversible domain where the lower bound of large deviations matches the upper bound;
(EL): Violation of the kinetic energy equality breaks the lower bound of large deviations;
(EU): The energy equality holds on a certain time-reversible closure of the weak-strong uniqueness regularity class;
(UN): The conditional equivalence of the quasi-potential and the Gaussian rate function is proved.
We explain why (UN) provides a potential connection between non-Gaussian large deviations in equilibrium and open problems of uniqueness for deterministic SQG equation. In general, the proof of a large deviation principle for invariant measures of stochastic PDEs relies on uniform dynamical large deviations, and the rate function is governed by the quasi-potential, which is defined as the infimum of the dynamical cost on the time interval $(-\infty,0]$ along paths in the space where the dynamical large deviations are proved. However, in this paper, the dynamical large deviation lower bound is shown to match the upper bound only when we restrict fluctuations to a closure of the so-called weak-strong uniqueness class $\mathcal{C}_{0}$ (see Definition 1.1 later on). This suggests that the quasi-potential should be defined in a restricted version, see (1.5), where the infimum is taken along paths corresponding to a subset $\mathcal{A}_{0}\subset\mathcal{C}_{0}$. To show that the quasi-potential (1.5) equals the Gaussian rate function $I_{Gauss}(\phi)=\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ for all $\phi\in H^{\alpha-2\beta}(\mathbb{T}^{2})$, a similar argument as in [BCF15] can be carried out. However, this relies on the existence of solutions for the deterministic SQG equation in $\mathcal{A}_{0}$. This links to the open problem of the uniqueness for the SQG equation.
1.1. Main results
Let $\alpha,\beta$ satisfy (1.2), and $T>0$ be a fixed time horizon in the whole context. We first state the result of dynamical large deviations for (1.3). Let
$\mathbb{X}_{\alpha,\beta}:=L^{2}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2}))\cap L^{2}_{w}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))\cap C([0,T];H^{\alpha-2\beta}_{w}(\mathbb{T}^{2})),$ (1.6)
where ”$w$” denotes the weak topology. Let $s>\alpha+1$, we regularize the noise by
$\xi_{\delta}=\sqrt{Q_{\delta}}\xi=(I+\delta\Lambda^{2s})^{-1/2}\xi.$
A detailed calculation concerning the blow-up speed of $\xi_{\delta}$ will be given in Section 2. For every $\theta\in\mathbb{X}_{\alpha,\beta}$, let
$\mathcal{I}_{dyna}(\theta)=\frac{1}{2}\bigg{\|}\partial_{t}\theta+\Lambda^{2\alpha}\theta+R^{\perp}\theta\cdot\nabla\theta\bigg{\|}^{2}_{L^{2}\left([0,T];H^{-2\beta}(\mathbb{T}^{2})\right)},\ \ \ \mathcal{I}_{0}(\theta(0))=\|\theta(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$
be large deviation costs for the dynamics and the initial data, respectively. Moreover, we set
$\mathcal{I}(\theta):=\mathcal{I}_{dyna}(\theta)+\mathcal{I}_{0}(\theta(0))$ (1.7)
to be the whole rate function. The following weak-strong uniqueness class and its closure are introduced as the spaces where we restrict the lower bound of large deviations.
Definition 1.1.
We say that $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ is a weak-strong uniqueness class of (1.1) for every $g\in L^{2}([0,T];L^{2}(\mathbb{T}^{2}))$ if, for arbitrary two weak solutions $\theta_{1},\theta_{2}$ (see Definition 4.1 later on) of the skeleton equation (1.1) with the same initial data $\theta_{1}(0)=\theta_{2}(0)$ and the same control $g$, we have $\theta_{1}=\theta_{2}$ in $\mathbb{X}_{\alpha,\beta}$ as long as $\theta_{1}\in\mathcal{C}_{0}$ and $\theta_{2}$ satisfies the $H^{\alpha-2\beta}$-energy inequality: for every $t\in[0,T],$
$\frac{1}{2}\|\theta_{2}(t)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\|\theta_{2}(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\left\|\theta_{2}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\langle\Lambda^{2\alpha-2\beta}\theta_{2},g\rangle\,\mathrm{d}s.$
Furthermore, the $\mathcal{I}$-closure of $\mathcal{C}_{0}$ is defined by
$\overline{\mathcal{C}_{0}}^{\mathcal{I}}:=\big{\{}\theta\in\mathbb{X}_{\alpha,\beta}:\text{there exists a sequence }\{\theta^{(n)}\}\subset\mathcal{C}_{0},\,\text{such that }\theta^{(n)}\rightarrow\theta\text{ in }\mathbb{X}_{\alpha,\beta},\,\mathcal{I}\big{(}\theta^{(n)}\big{)}\rightarrow\mathcal{I}(\theta)\big{\}}.$
In particular, examples of regularity classes $\mathcal{C}_{0}$ and $\overline{\mathcal{C}_{0}}^{\mathcal{I}}$ are provided in Appendix A. Furthermore, we provide the following assumptions of the initial data for (1.3).
Assumption A (Gaussian initial data).
Assume that the law of the initial data $\theta_{\varepsilon,\delta}(0)$ is $\mathcal{G}\left(0,\varepsilon Q_{\delta}/2\right)$, the Gaussian measure on $\dot{H}^{\alpha-2\beta}(\mathbb{T}^{2})$ with mean zero and covariance $\varepsilon Q_{\delta}/2.$
Remark 1.2.
Compared to the classical large deviations for stochastic PDEs, we allow for fluctuations in the initial data, which constitutes a theoretical generalization. In particular, for the case $\delta=0$, solutions of the Galerkin approximation equation of (1.3) starting from equilibrium exhibit the time-reversibility (see Lemma 7.2 below). We will later see how time-reversibility is related to the kinetic energy equality. We further remark that the above assumption can be extended to more general conditions (see [GHW24, Assumption 2.1]).
Based on the above preparation, the large deviations result is stated below.
Theorem 1.3 (Proposition 5.1, Proposition 6.3).
For every $\varepsilon,\delta>0$, let $\theta_{\varepsilon,\delta}$ be a stochastic generalized Leray solution of (1.3) in the sense of Definition 3.1 with initial data $\theta_{\varepsilon,\delta}(0)$ satisfying Assumption A. Let $\mu_{\varepsilon}=\mu_{\varepsilon,\delta}$ be the laws of $\theta_{\varepsilon,\delta}$ on $\mathbb{X}_{\alpha,\beta}$. Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ be a weak-strong uniqueness class of (1.1) and $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$. Assume that the scaling regime (2.7) holds for $(\varepsilon,\delta(\varepsilon))$. Then for any closed set $F\subset\mathbb{X}_{\alpha,\beta}$,
$\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon}(F)\leqslant-\inf_{\theta\in F}\mathcal{I}(\theta),$
and for any open set $G\subset\mathbb{X}_{\alpha,\beta}$,
$\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon}(G)\geqslant-\inf_{\theta\in G\cap\mathcal{C}}\mathcal{I}(\theta).$
We further comment that the above large deviations result holds for a sequence of Galerkin approximations of (1.3) (See (3.7) below) as well, see Proposition 5.1 and Proposition 6.3 for details. Based on this large deviations result, we provide a direct relationship between regularity classes of the weak-strong uniqueness and regularity classes where the energy equality holds. Concretely, we show that the kinetic energy equality holds on every time-reversible set of a certain closure of the weak-strong uniqueness class $\mathcal{C}_{0}$. This introduces the following result.
Theorem 1.4 (Theorem 7.1).
Assume that $\beta=\alpha/2$. Let $\mathcal{C}_{0}$ be a weak-strong uniqueness class of (1.1) and $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$. Let $\mathcal{R}\subset\mathcal{C}$ satisfy $\mathcal{R}=\mathfrak{T}_{T}\mathcal{R}$, where the time-reversal operator $\mathfrak{T}_{T}:\mathbb{X}_{\alpha,\alpha/2}\rightarrow\mathbb{X}_{\alpha,\alpha/2}$ is defined by $(\mathfrak{T}_{T}\theta)(\cdot):=-\theta(T-\cdot)$. Suppose that $\theta\in\mathcal{R}$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];L^{2}(\mathbb{T}^{2}))$ in the sense of Definition 4.1, then the following kinetic energy equality holds:
$\frac{1}{2}\|\theta(T)\|_{L^{2}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\|\theta(0)\|_{L^{2}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s.$
We point out that the objects of discussion in the above relationship are purely deterministic. However, the understanding is based on the perspective of probability and the theory of large deviations. Furthermore, the proof of Theorem 1.4 provides a sufficient condition for the validity of the energy equality (1.4) for arbitrary Leray solutions as well.
Sufficient condition: As long as one can prove a full large deviation principle for the Galerkin sequence of equation (1.3) with respect to the canonical rate function, as stated in Theorem 1.3, the deterministic energy equality will hold for arbitrary Leray solutions of (1.1).
More precisely, the following corollary implies that the violation of the kinetic energy equality breaks the lower bound of large deviations.
Corollary 1.5.
Let $\varepsilon>0$ and $m\in\mathbb{N}_{+}$ satisfy the scaling regime (2.8). Let $\theta_{\varepsilon,m}$ be a solution of (1.10) with initial data $\theta_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$ and let $\mu_{\varepsilon,m}$ be the law of $\theta_{\varepsilon,m}$ on $\mathbb{X}_{\alpha,\alpha/2}.$ Suppose that $\theta$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];L^{2}(\mathbb{T}^{2}))$ in the sense of Definition 4.1.
(i) Assume that
$\frac{1}{2}\|\theta(T)\|_{L^{2}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s>\frac{1}{2}\|\theta(0)\|_{L^{2}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s,$
then $\theta$ violates the large deviations lower bound:
$\inf_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2}:\theta\in G\}}\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)<-\mathcal{I}(\theta).$
(ii) Assume that
$\frac{1}{2}\|\theta(T)\|_{L^{2}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s<\frac{1}{2}\|\theta(0)\|_{L^{2}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s,$
then $\mathfrak{T}_{T}\theta$ violates the large deviations lower bound:
$\inf_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2},\,\mathfrak{T}_{T}\theta\in G\}}\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)<-\mathcal{I}(\mathfrak{T}_{T}\theta).$
Finally, we provide an analysis on the exact representation of the quasi-potential (1.5). Let $\mathcal{C}_{0}$ be a weak-strong uniqueness class on $[0,\infty)$ (See Definition 8.2). Denote
$\mathcal{A}_{0}:=\big{\{}\theta\in\mathcal{C}_{0}:\partial_{t}\theta\in L^{2}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2})),R^{\perp}\theta\cdot\nabla\theta\in L^{2}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2}))$ ,
[…]
Theorem 1.6 (Proposition 8.3).
Let $\mathcal{U}$ be the quasi-potential defined by (1.5). Then
(i) For every $\phi\in H^{\alpha-2\beta}(\mathbb{T}^{2})$, it holds that $\mathcal{U}(\phi)\geqslant\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$.
(ii) The equality $\mathcal{U}(\phi)=\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ holds for a given $\phi\in H^{\alpha-2\beta}(\mathbb{T}^{2})$ as long as the equation
$\left\{\begin{array}[]{l}\partial_{t}\bar{\theta}=-\Lambda^{2\alpha}\bar{\theta}-R^{\perp}\bar{\theta}\cdot\nabla\bar{\theta},\\ \bar{\theta}(0)=\phi,\end{array}\right.$
admits a unique weak solution in $\mathcal{A}_{0}$ in the sense of Definition 8.1.
1.2. Key idea and technical comment
Large deviations. Concerning the large deviations of (1.3), the main obstacles arise from two aspects. On the one hand, generalized Leray solutions of (1.3) are probabilistically weak and therefore cannot be represented as measurable maps of Brownian paths. This prevents us from using the well-known weak convergence approach by Budhiraja, Dupuis, and Maroulas [BDM08] for large deviations of stochastic PDEs. On the other hand, the uniqueness of the skeleton equation (1.1) is unknown in such critical and supercritical ($\alpha\in(0,1/2]$) regimes. Large deviations for subcritical stochastic SQG equations driven by multiplicative noise have been studied by Liu, Röckner and Zhu in [LRZ13], under the assumption that $\theta_{0}\in L^{p}\cap H^{\delta}$, where $\delta>2-2\alpha$ and $1/p\in(0,\alpha-1/2)$. In this case, the pathwise uniqueness of the stochastic SQG equations is obtained, and the skeleton equation also has a unique solution, which leads to the weak convergence approach. In general, the well-posedness of the skeleton equation plays a key role in large deviations. We refer readers to [Hey23] for more details, where a counterexample is proposed to illustrate the violation of the lower bound due to the lack of uniqueness. Also see [FG23, Section 8] for the consistency of the rate function and its lower semi-continuous envelope with the help of the uniqueness and stability of the skeleton equation. These two obstacles have also been encountered in the study of three-dimensional Landau-Lifshitz-Navier-Stokes equations and were resolved by Gess, Heydecker, and the second author [GHW24]. To summarize the key idea, the upper and lower bounds are investigated separately, with the lower bound being obtained manually by restricting to the weak-strong uniqueness class. In this context, we will adopt the same idea as in [GHW24] on the probabilistic side, but extend it on the analytic side. Specifically, we generalize this argument into an $\dot{H}^{-1/2}$-framework.
We consider a generalized Leray solution theory for (1.3) with an $\dot{H}^{-\gamma}$-energy inequality, where $\gamma=0$ or $1/2$. The negative Sobolev’s regularity directly produces obstacles in the well-definedness of the nonlinear term $u_{\theta_{\varepsilon}}\cdot\nabla\theta_{\varepsilon}$, since the pointwise product cannot be defined for distributions. Alternatively, a commutator approach is employed to make the nonlinear term well-defined. Specifically, for every $\theta,\phi\in C^{\infty}(\mathbb{T}^{2})$, we denote
$\left[\Lambda,\partial_{i}\phi\right](\Lambda^{-1}\theta):=\Lambda(\partial_{i}\phi\Lambda^{-1}\theta)-\partial_{i}\phi\theta$
as the commutator between $\Lambda$ and $\partial_{i}\phi$ for $i=1,2$. An intuitive computation shows the identity
$\int_{\mathbb{T}^{2}}\theta R^{\perp}\theta\cdot\nabla\phi\,\mathrm{d}x=\frac{1}{2}\int_{\mathbb{T}^{2}}R_{2}\theta\left[\Lambda,\partial_{1}\phi\right]\left(\Lambda^{-1}\theta\right)\,\mathrm{d}x-\frac{1}{2}\int_{\mathbb{T}^{2}}R_{1}\theta\left[\Lambda,\partial_{2}\phi\right]\left(\Lambda^{-1}\theta\right)\,\mathrm{d}x.$ (1.8)
Moreover, for every $\theta\in\mathbb{X}_{\alpha,\beta}$ and every $\phi\in C^{\infty}(\mathbb{T}^{2})$, the right-hand side of (1.8) is well-defined as well. This suggests that the right-hand side of (1.8) can be taken as the definition of the nonlinear term $\langle\nabla\cdot\left(\theta R^{\perp}\theta\right),\phi\rangle$, see Lemma 2.2 later on for more details.
Secondly, the compactness argument lays a fundamental in the study of both the existence of solutions and large deviations. In general, problems concerning the passage to the limit will be encountered in such arguments. Regarding the passage to the limit of the nonlinear term, the Littlewood-Paley theory is employed. Precisely, $\nabla\cdot\left(\theta R^{\perp}\theta\right)$ can be decomposed into three parts: for every $j\in\mathbb{Z}$,
$\nabla\cdot\left(\theta R^{\perp}\theta\right)=\nabla\cdot\left(\theta R^{\perp}\left(S_{j}\theta\right)\right)+\nabla\cdot\left(S_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right)+\nabla\cdot\left(H_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right),$ (1.9)
where $H_{j}\theta$ is the high-frequency part of $\theta$ and $S_{j}\theta$ is the low-frequency part defined by (2.1) in Section 2. It has been shown by Marchand in [Mar08a, Lemma 9.3] that the first two terms on the right-hand side of (1.9) are easier to estimate due to the regularizing effect of $S_{j}$. However, the estimation of $\nabla\cdot\left(H_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right)$ is more challenging and requires the commutator estimates. In particular, for every $\phi\in C^{\infty}(\mathbb{T}^{2})$, $\iota>0$ and $r\in(-1,1/2],$ we have
(${}_{\dot{H}^{-3-r-\iota}}\left\langle\nabla\cdot\left(H_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right),\phi\right\rangle_{\dot{H}^{3+r+\iota}}\leqslant C\|\phi\|_{\dot{H}^{3+r+\iota}(\mathbb{T}^{2})}\left\|H_{j}\theta\right\|_{\dot{H}^{-r}(\mathbb{T}^{2})}^{2}.$
The energy equality. In this part, we summarize the key idea of the proof for Theorem 1.4. The main idea arises from [GH23, Section 11] and [GHW24, Section 9]. In particular, in [GHW24, Section 9], the authors provide a purely analytic approach concerning the relationship between the energy equality and the weak-strong uniqueness class for three-dimensional forced Navier-Stokes equations. Moreover, they provide a brief explanation of its probabilistic interpretation. In this paper, we adopt the probabilistic approach for proving this relationship when $\beta=\alpha/2$.
The key ingredient is to utilize the fact that the Gaussian distribution $\mathcal{G}(0,\varepsilon P_{m}/2)$ is an invariant measure for the Galerkin approximation of (1.3):
$\partial_{t}\theta_{\varepsilon,m}=-\Lambda^{2\alpha}\theta_{\varepsilon,m}-P_{m}(u_{\theta_{\varepsilon,m}}\cdot\nabla\theta_{\varepsilon,m})+\varepsilon^{1/2}P_{m}\Lambda^{\alpha}\xi,$ (1.10)
where $P_{m}$ denotes the $m$-dimensional projection operator (see Section 3 for a rigorous definition). Thanks to the time-reversibility (See Lemma 7.2)
$\mathfrak{T}_{T}\theta_{\varepsilon,m}:=\left(-\theta_{\varepsilon,m}(T-t):0\leqslant t\leqslant T\right)\overset{\mathrm{d}}{=}\left(\theta_{\varepsilon,m}(t):0\leqslant t\leqslant T\right),$
both $\theta_{\varepsilon,m}$ and its time-reversal process $\mathfrak{T}_{T}\theta_{\varepsilon,m}$ satisfy large deviations as presented in Theorem 1.3. On the domain where the large deviations lower bound matches the upper bound ($\mathcal{C}$ in Theorem 1.3), the uniqueness of the large deviations rate function holds. Therefore, we have that $\mathcal{I}(\mathfrak{T}_{T}\theta)=\mathcal{I}(\theta)$ for all $\theta\in\mathbb{X}_{\alpha,\alpha/2}$ satisfying $\theta,\mathfrak{T}_{T}\theta\in\mathcal{C}$. For any weak solution $\theta$ of (1.1), direct computation shows that
$\mathcal{I}(\theta)=\|\theta(0)\|_{L^{2}(\mathbb{T}^{2})}^{2},\ \ \mathcal{I}(\mathfrak{T}_{T}\theta)=\|\theta(T)\|_{L^{2}(\mathbb{T}^{2})}^{2}+2\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s-2\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s.$
Then the kinetic energy equality holds as long as $\theta,\mathfrak{T}_{T}\theta\in\mathcal{C}$. Based on this observation, one can see that the uniqueness of the rate function plays a key role. Since the uniqueness holds on the domain where the lower bound matches the upper bound, then one can provide a sufficient condition concerning the open problem of the validity of the kinetic energy equality for arbitrary Leray solutions. As long as one can prove a full large deviation of (1.10) on $\mathbb{X}_{\alpha,\alpha/2}$ with respect to the rate function $\mathcal{I}$, then this implies that the kinetic energy equality holds on $\mathbb{X}_{\alpha,\alpha/2}$ without restriction.
1.3. Comments of literature
Large deviations for stochastic PDEs in singular limits. Large deviations of stochastic scalar conservation laws in the joint limits of vanishing noise and viscosity were studied by Mariani [Mar10]. In the work of Dirr, Fehrman and Gess [DFG24], large deviations for a conservative stochastic PDE corresponding to the simple symmetric exclusion process were established. Ferman and Gess [FG23] demonstrated that a class of generalized Dean-Kawasaki type equations satisfies large deviations with the same rate function as that of the zero range process. The second author and Zhang [WZ24] extended the analysis to a more general case concerning singular interactions for the Dean-Kawasaki equation. Gess, Heydecker, and the second author [GHW24] proved the large deviation principle for the Landau-Lifshitz-Navier-Stokes equations, whose rate function is consistent with the Quastel-Yau lattice gas model [QY98]. Further works concerning large deviations for stochastic PDEs in scaling limits with vanishing noise intensity and correlations can be found in [HW15, CD19a, CD19b, CP22].
Background of the energy functions. There are two types of energy that have physical significance for the SQG equation: the Hamiltonian and the kinetic energy. For the inviscid SQG equation, the functional
$\mathcal{H}(t):=\frac{1}{2}\|\theta(t)\|_{\dot{H}^{-1/2}\left(\mathbb{T}^{2}\right)}^{2}=\frac{1}{2}\int_{\mathbb{T}^{2}}\Lambda^{-1}\theta(t)\cdot\theta(t)\,\mathrm{d}x,\ \ t\in[0,T]$
can be taken as the Hamiltonian of the system. This is derived from the Euler-Poincaré variational principle. We refer readers to Arnold’s work [Arn66], which is an infinite-dimensional generalization of [Poi01]. (See [Res95, Section 2.2], [BSV19, Section 1.4, Appendix A.1] for more details.) Besides the Hamiltonian $\mathcal{H}$, the kinetic energy is widely discussed as well. In the system of the SQG equation, the field $\theta$ represents the temperature or surface buoyancy for a rapidly rotating stratified fluid with uniform potential vorticity, and $u_{\theta}=R^{\perp}\theta$ denotes the transport velocity field [HPGS95]. Since the Riesz transform is an $L^{2}$-isometry [SW71, Chapter VI, Section 2], the kinetic energy can be represented by
$\mathcal{K}(t)=\frac{1}{2}\int_{\mathbb{T}^{2}}|u_{\theta}(t)|^{2}\,\mathrm{d}x=\frac{1}{2}\int_{\mathbb{T}^{2}}|\theta(t)|^{2}\,\mathrm{d}x,\ \ t\in[0,T].$
The validity of the energy equality of the SQG equation with respect to both of these two kinds of energy has been widely investigated.
Energy equality for the SQG equation. Energy conservation of the inviscid SQG equation is related to Onsager’s conjecture [Ons49]. Isett and Vicol proved in [IV15] that $\theta\in L^{3}_{t}L^{3}_{x}$ implies the conservation of the Hamiltonian $\mathcal{H}$. For sufficiently regular solutions, the $L^{2}$-norm (the kinetic energy) is conserved as well. Several Besov-type regularity conditions are proposed for weak solutions to guarantee the conservation of the kinetic energy, see [Zho05], [Cha06a], [WYY23]. Dai [Dai17] investigated the kinetic energy equality for viscosity solutions of the super-critical dissipative SQG equation under some regularity conditions.
Weak-strong uniqueness for the SQG equation. Results of weak-strong uniqueness for the SQG equations with different fractional dissipation index $\alpha$ and different initial data $\theta_{0}$ can be found in the following literature. Constantin and Wu [CW99] proved the uniqueness of strong solution when $\alpha\in(1/2,1]$ with $\theta_{0}$ in $\dot{H}^{-1/2}$ or $L^{2}$ , which is analogous to the Ladyzhenskaya-Prodi-Serrin condition of Navier-Stokes equation [KL57, Pro59, Ser62]. In the critical and supercritical cases $\alpha\in(0,1/2]$, there are several results for $L^{2}$-valued initial data. Dong and Chen in [DC06] proposed Ladyzhenskaya-Prodi-Serrin type condition $L^{r}_{t}L^{p}_{x}$ to $\nabla\theta$ and showed a weak-strong uniqueness property. In [DC12], the authors extended this regularity condition to Besov space $L^{r}_{t}B^{0}_{p,\infty}$. For the critical case $\alpha=1/2$, Marchand [Mar08b] proved the weak-strong uniqueness in BMO-type space. We will summarize examples of weak-strong uniqueness for subcritical, critical, and supercritical cases in Appendix A, and illustrate that these regularity classes can be taken as examples of the domain where we restrict the lower bound of large deviations. Furthermore, we point out that, to the best of our knowledge, weak-strong uniqueness results for the supercritical case $\alpha\in(0,1/2)$ with $\theta_{0}$ in $\dot{H}^{-1/2}$ have not been obtained. In this case, the dissipation term is not sufficient to prevent weak solutions from being distribution-valued. Therefore, it is challenging to propose a reasonable regularity condition for proving the uniqueness.
Relationship to the study of convex integration. From the previous explanations, it can be inferred that obtaining matching upper and lower bounds for large deviations is related to the uniqueness of solutions for (1.1). Recently, the convex integration technique has led to various non-uniqueness results for fluid equations. Buckmaster, Shkoller, and Vicol [BSV19] demonstrated the non-uniqueness of weak solutions for both inviscid and dissipative SQG equations by converting the equation into momentum form. Subsequently, different convex integration schemes have been employed to construct non-unique solutions for inviscid and dissipative SQG equations, as illustrated in [CKL21] and [IM21]. All of these works focus on the SQG equations without external forces, where the solutions satisfy $\Lambda^{-1}\theta\in C_{t}^{\sigma}C_{x}^{s}$ for some $\sigma$ and $s$. However, it remains an open question whether $\theta\in L^{2}_{t}H^{\alpha}_{x}$, which is the space considered in Theorem 1.4. For example, in [BSV19], the solutions were constructed in $C_{t}^{\sigma}C_{x}^{s}$ with $51/2<s<4/$ 5 and $\sigma<s/(2-s).$ Therefore, these non-uniqueness results do not imply a violation of the full large deviations.
In the context of stochastic SQG equations, Yamazaki established the non-uniqueness of solutions driven by trace-class noise [Yam22], and later extended these results to include linear multiplicative noise [Yam23]. In [HZZ23] and [HLZZ24], the authors discussed the cases of irregular spatial noise and space-time white noise, respectively. We remark that, in [HLZZ24], infinitely many non-Gaussian ergodic stationary solutions of (1.3) (with $\delta=0$) in $B^{-1/2}_{p,1},p\geqslant 2,$ were constructed. We hope to connect the discussion of non-Gaussian stationary solutions with the uniqueness problem of deterministic SQG equations through the non-Gaussian large deviations in Section 8.
1.4. Structure of the paper
This paper is organized as follows. In Section 2, we introduce some preliminaries which will be used throughout the paper. Section 3 is dedicated to proving the existence and exponential tightness for stochastic generalized Leray solutions. Section 4 presents a variational characterization of the rate function. In Section 5, we establish the upper bound for large deviations. Section 6 provides a restricted lower bound for large deviations using the entropy method. In Section 7, we prove a relationship between the weak-strong uniqueness class and the energy equality for (1.1). Finally, in Section 8, we discuss an explicit representation of the quasi-potential. Additionally, Appendix A summarizes several examples of weak-strong uniqueness classes concerning different ranges of parameters $\alpha$ and $\beta$, and Appendix B demonstrates the large deviations for Gaussian initial conditions.
2. Preliminary
This section is devoted to presenting some preliminaries.
2.1. Notations and spaces
Throughout the paper, we use the notation $a\lesssim b$ to denote that there exists a constant $C>0$ such that $a\leqslant Cb$, where $C$ may change from line to line.
By testing the constant function $1$ for (1.1), one can see that the spatial average of the solutions is conserved. Therefore, we assume that the zero Fourier mode of the initial data is zero. In this way, we study (1.1) and (1.3) in spaces with zero mean.
For any two topological spaces $X_{1}$ and $X_{2}$, $C(X_{1};X_{2})$ is defined as the space of continuous functions from $X_{1}$ to $X_{2}$. $C_{b}(X_{1})$ is defined as the space of bounded continuous functions from $X_{1}$ to $\mathbb{R}$. Given a Banach space $E$ with norm $\|\cdot\|_{E}$, for any $\sigma\in(0,1)$, we write $C^{\sigma}([0,T];E)$ to denote the space of $\sigma$-Hölder continuous functions from $[0,T]$ to $E$ endowed with the semi-norm
$\|f\|_{C^{\sigma}([0,T];E)}:=\sup_{s,t\in[0,T],s\neq t}\frac{\|f(s)-f(t)\|_{E}}{|t-s|^{\sigma}}.$
Given $p>1,\sigma\in(0,1)$, let $W^{\sigma,p}([0,T];E)$ denote the space of all $f\in L^{p}([0,T];E)$ such that
$\int_{0}^{T}\int_{0}^{T}\frac{\|f(t)-f(s)\|_{E}^{p}}{|t-s|^{1+\sigma p}}\mathrm{\,d}t\,\mathrm{d}s<\infty,$
endowed with the norm
$\|f\|_{W^{\sigma,p}([0,T];E)}^{p}:=\int_{0}^{T}\|f(t)\|_{E}^{p}\mathrm{\,d}t+\int_{0}^{T}\int_{0}^{T}\frac{\|f(t)-f(s)\|_{E}^{p}}{|t-s|^{1+\sigma p}}\mathrm{\,d}t\,\mathrm{d}s.$
We denote the distribution space on $\mathbb{T}^{2}$ by $\mathscr{D}^{\prime}(\mathbb{T}^{2})$, which is the dual space of $C^{\infty}(\mathbb{T}^{2})$. Let $f\in\mathscr{D}^{\prime}(\mathbb{T}^{2})$ and let $g=(g(k))_{k\in\mathbb{Z}^{2}}$ be at most of polynomial growth. The Fourier transform of $f$ and the inverse Fourier transform of $g$ are defined as
$\mathcal{F}_{\mathbb{T}^{2}}f(k):=\hat{f}(k):=\int_{\mathbb{T}^{2}}e^{2\pi\mathrm{i}x\cdot k}f(x)\mathrm{\,d}x,\quad k\in\mathbb{Z}^{2},\quad\mathcal{F}_{\mathbb{T}^{2}}^{-1}g(x):=\sum_{k\in\mathbb{Z}^{2}}e^{2\pi\mathrm{i}x\cdot k}g(k),\quad x\in\mathbb{T}^{2}.$
[…]
$\left\{\sqrt{2}\sin(2\pi k\cdot)|k\in\mathbb{Z}_{+}^{2}\right\}\cup\left\{\sqrt{2}\cos(2\pi k\cdot)|k\in\mathbb{Z}_{-}^{2}\right\}$
is an orthonormal eigenbasis of $\Lambda$ on $H$, which we denote by $\{e_{k}\}$. For every $s>0$, define
$\|f\|_{H^{s}(\mathbb{T}^{2})}^{2}:=\sum_{k\in\mathbb{Z}^{2}}|2\pi k|^{2s}\big{|}\hat{f}(k)\big{|}^{2},$
and let $H^{s}(\mathbb{T}^{2})$ denote the space of all $f\in H$ for which $\|f\|_{H^{s}(\mathbb{T}^{2})}$ is finite. Define $H^{-s}(\mathbb{T}^{2})$ to be the dual of $H^{s}(\mathbb{T}^{2})$. Finally, let $H^{\infty}(\mathbb{T}^{2}):=\bigcap\limits_{s>0}H^{s}(\mathbb{T}^{2})$.
Now we recall the Littlewood-Paley decomposition briefly. Let $\varphi\in C_{c}^{\infty}\left(\mathbb{R}^{2}\right)$ be a non-negative radial function so that $\varphi(\xi)=1$ for $|\xi|\leqslant 1/2$ and $\varphi(\xi)=0$ for $|\xi|\geqslant 1$. Let $\psi(\xi)=\varphi(\xi/2)-\varphi(\xi)$. Then $\big{(}\psi(\cdot/2^{j})\big{)}_{j\in\mathbb{Z}}$ is a dyadic partition of unity. For every $j\in\mathbb{Z}$, we define the $j$-th dyadic block of the Littlewood-Paley decomposition of $f\in\mathscr{D}^{\prime}(\mathbb{T}^{2})$ by
$\Delta_{j}f:=\mathcal{F}_{\mathbb{T}^{2}}^{-1}\left(\psi\left(\cdot/2^{j}\right)\hat{f}\right).$
The low-frequencies and high-frequencies cutting-off operators are defined by
$S_{j}f:=\mathcal{F}_{\mathbb{T}^{2}}^{-1}\left(\varphi\left(\cdot/2^{j}\right)\hat{f}\right)=\sum_{k=-\infty}^{j-1}\Delta_{k}f,$ (2.1)
$H_{j}f:=\left(Id-S_{j}\right)f=\sum_{k=j}^{\infty}\Delta_{k}f.$
For every $\phi\in H^{\infty}(\mathbb{T}^{2})$, the commutator between $\Lambda$ and $\phi$ is defined by $[\Lambda,\phi]g:=\Lambda(\phi g)-\phi\Lambda g$, for any $g\in H^{\infty}(\mathbb{T}^{2}).$
The following lemma enables us to introduce the commutator estimate, which will contribute to the well-definedness of the nonlinear term.
Lemma 2.1.
Let $\theta,\phi\in H^{\infty}\left(\mathbb{T}^{2}\right)$, then
$2\int_{\mathbb{T}^{2}}\theta R^{\perp}\theta\cdot\nabla\phi\,\mathrm{d}x=\int_{\mathbb{T}^{2}}R_{2}\theta\left[\Lambda,\partial_{1}\phi\right]\left(\Lambda^{-1}\theta\right)\,\mathrm{d}x-\int_{\mathbb{T}^{2}}R_{1}\theta\left[\Lambda,\partial_{2}\phi\right]\left(\Lambda^{-1}\theta\right)\,\mathrm{d}x.$
Lemma 2.1 can be proved by integration by parts, see for example, [Mar08a, Lemma 2.1].
Lemma 2.2 (Commutator Estimate).
For every $r\in(-1,1)$, let $\phi\in H^{\infty}\left(\mathbb{T}^{2}\right)$ and $g\in H^{r}\left(\mathbb{T}^{2}\right)$, then
$\|[\Lambda,\phi]g\|_{H^{r}(\mathbb{T}^{2})}\leqslant C(r,\phi)\|g\|_{H^{r}(\mathbb{T}^{2})}.$
This commutator estimate is stated for $\mathbb{R}^{2}$ functions in [Mar08a, Lemma 2.2]. The same proof also works in the case of $\mathbb{T}^{2}$ functions, see [ZZ14, Lemma 4.2], [BSV19, Lemma A.5].
It follows from the above two lemmas that for every $f$ in $H^{-1/2}\left(\mathbb{T}^{2}\right)$, $\nabla\cdot\left(fR^{\perp}(f)\right)$ can be defined as a distribution by
$\nabla\cdot\left(fR^{\perp}(f)\right)=\nabla\cdot\left(fR^{\perp}\left(S_{j}f\right)\right)+\nabla\cdot\left(S_{j}fR^{\perp}\left(H_{j}f\right)\right)+\nabla\cdot\left(H_{j}fR^{\perp}\left(H_{j}f\right)\right),$ (2.2)
for arbitrary $j\in\mathbb{Z}$. And $\nabla\cdot\left(H_{j}fR^{\perp}\left(H_{j}f\right)\right)$ is the distribution defined by
$\left\langle\nabla\cdot\left(H_{j}fR^{\perp}\left(H_{j}f\right)\right),\phi\right\rangle:=$ $-{}_{H^{-1/2}}\big{\langle}\frac{1}{2}R_{2}\left(H_{j}f\right),\left[\Lambda,\partial_{1}\phi\right]\Lambda^{-1}\left(H_{j}f\right)\big{\rangle}_{H^{1/2}}$ (2.3)
$+{}_{H^{-1/2}}\big{\langle}\frac{1}{2}R_{1}\left(H_{j}f\right),\left[\Lambda,\partial_{2}\phi\right]\Lambda^{-1}\left(H_{j}f\right)\big{\rangle}_{H^{1/2}}$
for all $\phi\in H^{\infty}\left(\mathbb{T}^{2}\right)$. More precisely, $\nabla\cdot\left(fR^{\perp}\left(S_{j}f\right)\right)$ and $\nabla\cdot\left(S_{j}fR^{\perp}\left(H_{j}f\right)\right)$ are well-defined since $S_{j}f\in L^{\infty}(\mathbb{T}^{2})$. For the high-frequency part $\nabla\cdot\left(H_{j}fR^{\perp}\left(H_{j}f\right)\right)$, it follows from Lemma 2.2 that, for any $r\in(-1,1/2]$, $j\in\mathbb{Z}$,
$\left|\langle\nabla\cdot(H_{j}fR^{\perp}(H_{j}f)),\phi\rangle\right|\leqslant C(r,\phi)\|H_{j}f\|_{H^{-r}(\mathbb{T}^{2})}\|H_{j}f\|_{H^{r-1}(\mathbb{T}^{2})}\leqslant C(r,\phi)\|H_{j}f\|_{H^{-r}(\mathbb{T}^{2})}^{2}.$ (2.4)
Hence (2.3) is well-defined. More concretely, $C(r,\phi)$ can be controlled by $C(\iota)\|\phi\|_{H^{3+r+\iota}(\mathbb{T}^{2})}$ for some $\iota>0.$ The definition (2.3) is independent of the choice of $j\in\mathbb{Z}$.
2.2. Basic results of large deviations
We recall some basic concepts and results in the theory of large deviations. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and let $X$ be a topological space with the Borel $\sigma$-field $\mathcal{B}(X)$.
Definition 2.3 (Rate function).
Let $I$ be a function from $X$ to $[0,\infty]$. We say that $I$ is a rate function on $X$ if $I$ is lower semi-continuous on $X$.
Definition 2.4 (Large Derivations).
Let $I$ be a rate function on $X$. Suppose that $\left\{\mu_{\varepsilon}\right\}_{\varepsilon>0}$ is a family of probability measures on $X$. We say that $\left\{\mu_{\varepsilon}\right\}_{\varepsilon>0}$ satisfies the large deviations on $X$ with speed $\varepsilon^{-1}$ and rate function $I$, if
(i) (Upper bound) For any closed subset $F\subset X$,
$\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon}(F)\leqslant-\inf_{x\in F}I(x).$
(ii) (Lower bound) For any open subset $G\subset X$,
$\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon}(G)\geqslant-\inf_{x\in G}I(x).$
The following entropy method will be used in our proof of lower bound.
Lemma 2.5.
[Mar10, Lemma 7],[GHW24, Lemma 3.2] Let $E$ be a separable and Hausdorff space, $I$ be a function from $E$ to $[0,+\infty],$ and $\{\mu_{\varepsilon}\}_{\varepsilon>0}$ be a family of probability measures on $E$. Then $\{\mu_{\varepsilon}\}$ satisfies the large deviation lower bound with speed $\varepsilon^{-1}$ and rate function $I$ if and only if, for every $x\in E$, there exists a sequence of probability measures $\{\pi_{\varepsilon,x}\}_{\varepsilon>0}$ that satisfies the following two conditions.
(i) (Weak convergence) $\pi_{\varepsilon,x}\rightarrow\delta_{x}$ weakly, as $\varepsilon\rightarrow 0$.
(ii) (The entropy inequality)
$\limsup_{\varepsilon\rightarrow 0}\varepsilon{\rm{Ent}}(\pi_{\varepsilon,x}|\mu_{\varepsilon})\leqslant I(x).$ (2.5)
2.3. The Regularization of Noise
Recall that $\{e_{k}\}_{k\in\mathbb{Z}^{2}\backslash\{0\}}$ is an orthonormal basis of $H$. A cylindrical Wiener process $W$ on $H$ has the following representation,
$W(t)=\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}\beta_{k}(t)e_{k},\quad t\in[0,T],$
where $\{\beta_{k}\}_{k\in\mathbb{Z}^{2}\backslash\{0\}}$ is a family of independent real-valued Brownian motions. For every $\delta>0$ and fixed $s>\alpha+1$, we define the regularizing operator $Q_{\delta}$ as
$Q_{\delta}=(I+\delta(-\Delta)^{s})^{-1},$
then
$\sqrt{Q_{\delta}}e_{k}=(I+\delta|k|^{2s})^{-1/2}e_{k}=:\lambda_{\delta,k}e_{k},\quad k\in\mathbb{Z}^{2}\backslash\{0\}.$
For a bounded operator $T:H\rightarrow H$, the Hilbert-Schmidt norm of $T$ is defined as
$\|T\|_{HS}^{2}:=\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}\|Te_{k}\|_{H}^{2}.$ (2.6)
A direct computation shows that
$\left\|\Lambda^{\alpha}\circ\sqrt{Q_{\delta}}\right\|_{HS}^{2}$ $=\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}|k|^{2\alpha}\lambda_{\delta,k}^{2}=\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}\frac{|k|^{2\alpha}}{1+\delta|k|^{2s}}$
$\lesssim\int_{1}^{\infty}\frac{r^{2\alpha+1}}{1+\delta r^{2s}}\,\mathrm{d}r\lesssim\delta^{-\frac{\alpha+1}{s}}\int_{\delta^{\frac{1}{2s}}}^{\infty}\frac{u^{2\alpha+1}}{1+u^{2s}}\,\mathrm{d}u\lesssim\delta^{-\frac{\alpha+1}{s}}.$
This indicates the blow-up speed of solutions of (1.3) and is therefore related to the scaling regimes we will encounter later on.
In the study of large deviations, we restrict that the scaling regime $(\varepsilon,\delta(\varepsilon))$ satisfies
$\lim_{\varepsilon\rightarrow 0}\varepsilon\delta(\varepsilon)^{-\frac{\alpha+1}{s}}=0.$ (2.7)
Similarly, we require that the scaling regime $(\varepsilon,m(\varepsilon))$ for the Galerkin sequence (3.7) satisfies
$\lim_{\varepsilon\rightarrow 0}\varepsilon m(\varepsilon)^{2+2\alpha}=0.$ (2.8)
In this context, we denote $\xi=\mathrm{d}W(t)/\mathrm{d}t$ and $\xi_{\delta}=\sqrt{Q_{\delta}}\mathrm{d}W(t)/\mathrm{d}t.$
3. Existence and Exponential Tightness
In this section, we will establish the existence of solutions of the regularized stochastic PDE (1.3) for every $\varepsilon,\delta>0$, and the exponential tightness under the scaling regimes (2.7) and (2.8).
First, we provide the definition of solutions. The following definition can be regarded as a generalization of the Leray solution. Since the Sobolev exponent is negative when $\beta=\alpha/2+1/4$, a Littlewood-Paley analysis is needed to make the nonlinear term well-defined.
Definition 3.1.
We say that $(\theta_{\varepsilon,\delta},W)$, $\left(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\in[0,T]},\mathbb{P}\right)$ is a stochastic generalized Leray solution of (1.3) with initial data $\theta_{0}$ if, $\theta_{0}$ is an $H^{\alpha-2\beta}(\mathbb{T}^{2})$-valued random element, $W$ is a cylindrical Wiener process on $H^{\alpha-2\beta}(\mathbb{T}^{2})$ independent of $\theta_{0}$, and $\theta_{\varepsilon,\delta}$ is a progressively measurable process satisfying:
(i) $\mathbb{P}$-almost surely, $\theta_{\varepsilon,\delta}\in L^{\infty}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2}))\cap L^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))\cap C([0,T];H_{w}^{\alpha-2\beta}(\mathbb{T}^{2}))$.
(ii) $\mathbb{P}$-almost surely, for every $\varphi\in C^{\infty}\left([0,T];H^{\infty}(\mathbb{T}^{2})\right)$, for all $t\in[0,T]$,
$\left\langle\theta_{\varepsilon,\delta}(t),\varphi(t)\right\rangle+\int_{0}^{t}\left\langle\theta_{\varepsilon,\delta}(s),\Lambda^{2\alpha}\varphi(s)\right\rangle\,\mathrm{d}s=\left\langle\theta_{0},\varphi(0)\right\rangle+\int_{0}^{t}\langle\theta_{\varepsilon,\delta}(s),\partial_{s}\varphi(s)\rangle\,\mathrm{d}s$
$-\frac{1}{2}\int_{0}^{t}\left\langle R_{2}\theta_{\varepsilon,\delta}(s),\left[\Lambda,\partial_{1}\varphi(s)\right]\Lambda^{-1}\theta_{\varepsilon,\delta}(s)\right\rangle\,\mathrm{d}s+\frac{1}{2}\int_{0}^{t}\left\langle R_{1}\theta_{\varepsilon,\delta}(s),\left[\Lambda,\partial_{2}\varphi(s)\right]\Lambda^{-1}\theta_{\varepsilon,\delta}(s)\right\rangle\,\mathrm{d}s$
$+\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\beta}\varphi(s),\sqrt{Q_{\delta}}\,\mathrm{d}W(s)\rangle.$
Here and subsequently, the nonlinear terms $\left\langle R_{j}\theta_{\varepsilon,\delta}(s),\left[\Lambda,\partial_{i}\varphi(s)\right]\Lambda^{-1}\theta_{\varepsilon,\delta}(s)\right\rangle(i,j=1,2)$ are defined by the $H^{-1/2}-H^{1/2}$ dual as in (2.3).
(iii) $\theta_{\varepsilon,\delta}$ satisfies the pathwise $H^{\alpha-2\beta}$-energy inequality: $\mathbb{P}$-almost surely, for all $t\in[0,T]$,
$\frac{1}{2}\left\|\theta_{\varepsilon,\delta}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+$ $\int_{0}^{t}\left\|\theta_{\varepsilon,\delta}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\left\|\theta_{0}\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$
$+\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,\delta}(s),\sqrt{Q_{\delta}}\,\mathrm{d}W(s)\rangle+\frac{\varepsilon}{2}\left\|\Lambda^{\alpha}\circ\sqrt{Q_{\delta}}\right\|_{HS}^{2}t.$ (3.1)
Theorem 3.2 (Existence of stochastic generalized Leray solutions).
[…]
Proof.
The existence of a martingale solution is established in [ZZ14, Theorem 4.5], following from a similar argument in [FG95]. And the pathwise $H^{\alpha-2\beta}$-energy inequality can be derived in a similar way as in the proof of [GHW24, Proposition 5.1]. We only sketch the proof for the pathwise $H^{\alpha-2\beta}$-energy inequality.
We denote $H_{m}:=\mathrm{span}\{e_{k}\}_{0<|k|\leqslant m}.$ Let $P_{m}$ be the projection operator from $H$ to $H_{m}$. Now we consider the Galerkin approximation equations
$\mathrm{d}\tilde{\theta}_{\varepsilon,\delta,m}(t)=-\Lambda^{2\alpha}\tilde{\theta}_{\varepsilon,\delta,m}(t)\,\mathrm{d}t-P_{m}\left(u_{\tilde{\theta}_{\varepsilon,\delta,m}}(t)\cdot\nabla\tilde{\theta}_{\varepsilon,\delta,m}(t)\right)\,\mathrm{d}t+\varepsilon^{1/2}\Lambda^{2\beta}P_{m}\sqrt{Q_{\delta}}\,\mathrm{d}\tilde{W}(t),$ (3.2)
$\tilde{\theta}_{\varepsilon,\delta,m}(0)=P_{m}\tilde{\theta}_{0}.$
Since the coefficients in (3.2) satisfy the local Lipschitz condition, there exists a local solution on $[0,\tau_{K}]$ for any $K>0$. Here the stopping time $\tau_{K}$ is defined by $\tau_{K}:=\inf\{t\in[0,T]:\|\tilde{\theta}_{\varepsilon,\delta,m}(t)\|_{H}^{2}>K\}\wedge T$ for any $t\in[0,T].$ By the finite dimensional Itô’s formula, $\tilde{\mathbb{P}}$-almost surely, for all $t\in[0,\tau_{K})$,
$\frac{1}{2}\left\|\tilde{\theta}_{\varepsilon,\delta,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ $+\int_{0}^{t}\left\|\tilde{\theta}_{\varepsilon,\delta,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\left\|P_{m}\tilde{\theta}_{0}\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ (3.3)
$+\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\alpha-2\beta}\tilde{\theta}_{\varepsilon,\delta,m}(s),P_{m}\sqrt{Q_{\delta}}\,\mathrm{d}\tilde{W}(s)\rangle+\frac{\varepsilon}{2}\left\|\Lambda^{\alpha}\circ P_{m}\sqrt{Q_{\delta}}\right\|_{HS}^{2}t.$
Indeed, in the case $\beta=\alpha/2$, since $\nabla\cdot u_{\tilde{\theta}_{\varepsilon,\delta,m}}=0$, it holds that
$\langle u_{\tilde{\theta}_{\varepsilon,\delta,m}}\cdot\nabla\tilde{\theta}_{\varepsilon,\delta,m},\tilde{\theta}_{\varepsilon,\delta,m}\rangle=-\langle u_{\tilde{\theta}_{\varepsilon,\delta,m}}\tilde{\theta}_{\varepsilon,\delta,m},\nabla\tilde{\theta}_{\varepsilon,\delta,m}\rangle=0.$
And in the case $\beta=\alpha/2+1/4$, it holds that
$\langle u_{\tilde{\theta}_{\varepsilon,\delta,m}}\cdot\nabla\tilde{\theta}_{\varepsilon,\delta,m},\Lambda^{-1}\tilde{\theta}_{\varepsilon,\delta,m}\rangle=-\langle u_{\tilde{\theta}_{\varepsilon,\delta,m}}\tilde{\theta}_{\varepsilon,\delta,m},\nabla\Lambda^{-1}\tilde{\theta}_{\varepsilon,\delta,m}\rangle=-\langle\tilde{\theta}_{\varepsilon,\delta,m}R^{\perp}\tilde{\theta}_{\varepsilon,\delta,m},R\tilde{\theta}_{\varepsilon,\delta,m}\rangle=0.$
Therefore, (3.3) holds. Using Burkholder-Davis-Gundy inequality, we get the estimate
$\mathbb{\tilde{E}}\sup_{t\in[0,\tau_{K})}\|\tilde{\theta}_{\varepsilon,\delta,m}(t)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+2\mathbb{\tilde{E}}\int_{0}^{\tau_{K}}\|\tilde{\theta}_{\varepsilon,\delta,m}(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant C(T),$ (3.4)
[…]
$\mathbb{E}$ $\left[\int_{0}^{T}\psi(t)\left(\frac{1}{2}\left\|\theta_{\varepsilon,\delta,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}-\frac{1}{2}\left\|\theta_{\varepsilon,\delta,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\theta_{\varepsilon,\delta,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right.\right.$ (3.5)
$+\left.\left.\sqrt{\varepsilon}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\,\theta_{\varepsilon,\delta,m}(s),P_{m}\sqrt{Q_{\delta}}\mathrm{d}W(s)\right\rangle-\frac{\varepsilon}{2}\left\|\Lambda^{\alpha}\circ P_{m}\sqrt{Q_{\delta}}\right\|_{HS}^{2}t\right)\,\mathrm{d}t\right]=0.$
Set $M^{m}_{1}(t):=\int_{0}^{t}\sqrt{\varepsilon}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,\delta,m}(s),P_{m}\sqrt{Q_{\delta}}\,\mathrm{d}W(s)\right\rangle$, then $\{M^{m}_{1}\}_{m\geqslant 1}$ is tight in $C([0,T];\mathbb{R}).$ We denote the limit by $M_{1}:$
$\mathbb{E}\left[\int_{0}^{T}\psi(t)M_{1}(t)\,\mathrm{d}t\right]=\lim_{m\rightarrow\infty}\mathbb{E}\left[\int_{0}^{T}\psi(t)M_{1}^{m}(t)\,\mathrm{d}t\right].$
Letting $m\rightarrow\infty$ in (3.5), it follows from the lower semi-continuity of the Sobolev norms and Fatou’s lemma that
$\mathbb{E}\left[\int_{0}^{T}\psi(t)\left(\frac{1}{2}\left\|\theta_{\varepsilon,\delta}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right.\right.$ $-\frac{1}{2}\left\|\theta_{\varepsilon,\delta}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$
$+\int_{0}^{t}\left\|\theta_{\varepsilon,\delta}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$ $\left.\left.+M_{1}(t)-\frac{\varepsilon}{2}\left\|\Lambda^{\alpha}\circ P_{m}\sqrt{Q_{\delta}}\right\|_{HS}^{2}t\right)\,\mathrm{d}t\right]\leqslant 0.$
Since $\psi\in L^{\infty}\left(\mathbb{P};L^{\infty}\left([0,T],\mathbb{R}_{+}\right)\right)$ is arbitrary, $\mathbb{P}$-almost surely, for all $t\in[0,T],$
$\frac{1}{2}\left\|\theta_{\varepsilon,\delta}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\leqslant\frac{1}{2}\left\|\theta_{\varepsilon,\delta}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}-\int_{0}^{t}\left\|\theta_{\varepsilon,\delta}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s-M_{1}(t)+\frac{\varepsilon}{2}\left\|\Lambda^{\alpha}\circ\sqrt{Q_{\delta}}\right\|_{HS}^{2}t.$
To establish the pathwise $H^{\alpha-2\beta}$-energy inequality (3.1), it is sufficient to verify that
$M_{1}(t)=\sqrt{\varepsilon}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,\delta}(s),\sqrt{Q_{\delta}}\,\mathrm{d}W(s)\right\rangle.$ (3.6)
A detailed proof of (3.6) can be found in [GHW24, Proposition 5.1]. ∎
If we set $\delta=0$ and $\theta_{m,0}=P_{m}\theta_{0}\in H_{m}$, then (3.2) becomes
$\mathrm{d}\theta_{\varepsilon,m}(t)=-\Lambda^{2\alpha}\theta_{\varepsilon,m}\,\mathrm{d}t-P_{m}\left(u_{\theta_{\varepsilon,m}}\cdot\nabla\theta_{\varepsilon,m}\right)\,\mathrm{d}t+\varepsilon^{1/2}\Lambda^{2\beta}P_{m}\,\mathrm{d}W(t),$ (3.7)
$\theta_{\varepsilon,m}(0)=\theta_{m,0}.$
For any $\varepsilon>0$ and $m\in\mathbb{N}_{+}$, (3.7) admits a solution $\theta_{\varepsilon,m}$ as well. For any $\varepsilon,\delta>0$, let $\theta_{\varepsilon,\delta}$ be a stochastic generalized Leray solution of (1.3). We denote $\mu_{\varepsilon,\delta}$ and $\mu_{\varepsilon,m}$ as the laws of $\theta_{\varepsilon,\delta}$ and $\theta_{\varepsilon,m}$, respectively. In the remainder of this section, we will show that $\{\mu_{\varepsilon,\delta}\}$ and $\{\mu_{\varepsilon,m}\}$ are exponentially tight in $\mathbb{X}_{\alpha,\beta}$ under (2.7) and (2.8) respectively. To prove this, we require some exponential estimate lemmas.
The following lemma is aimed at obtaining an estimate in the space $L^{\infty}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2}))\cap L^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))$.
Lemma 3.3.
Let $\varepsilon>0$ and $m\in\mathbb{N}$. Suppose that $\theta_{\varepsilon,m}$ is a solution of (3.7) with $\theta_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$, then there exists $\eta_{0}>0$ such that for every $\eta\in\left[0,\eta_{0}\right]$, under the scaling regime (2.8), we have
$\varepsilon\log\mathbb{E}\exp\left\{\frac{\eta}{2\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{2\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\leqslant C\left(\eta,T\right),$
and
$\limsup_{R\rightarrow\infty}\sup_{\varepsilon\in(0,1)}\varepsilon\log\mathbb{P}\left(\frac{\eta}{2\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{2\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\geqslant R\right)=-\infty.$
Proof.
[…]
$\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ $=\left\|\theta_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}-\int_{0}^{t}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$ (3.8)
$+\varepsilon^{1/2}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle+\frac{\varepsilon}{2}t\left\|P_{m}\Lambda^{\alpha}\right\|_{HS}^{2}.$
It follows from (3.8), Burkholder-Davis-Gundy inequality and the scaling regime (2.8) that
$\mathbb{E}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\mathbb{E}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\mathbb{E}\left\|\theta_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+CT.$ (3.9)
This implies that $\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle$ has bounded quadratic variation, and therefore it is a martingale. Consequently, for any $\eta>0$, the process
$\exp\left\{-\frac{\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}$
is a submartingale. By Doob’s maximal inequality, we see that
$\mathbb{E}\left[\sup_{t\in[0,T]}\right.$ $\left.\exp\left\{-\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right]$ (3.10)
$\leqslant 4\mathbb{E}\left[\exp\left\{-\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right].$
Taking the exponential function of both sides of (3.8) and using (2.8) yield that
$\exp\left\{\frac{\eta}{\varepsilon}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right\}\cdot\exp\left\{\frac{\eta}{\varepsilon}\int_{0}^{t}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$ (3.11)
$\leqslant$ $\exp\big{\{}\frac{C(\eta)}{\varepsilon}T\big{\}}\cdot\exp\left\{\frac{\eta}{\varepsilon}\left\|\theta_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right\}\cdot\exp\left\{-\frac{\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}.$
For any $\eta\in(0,1/2)$, according to the scaling regime (2.8), the initial data part can be calculated by
$\quad\varepsilon\log\mathbb{E}\Big{(}\exp\Big{\{}\frac{\eta}{\varepsilon}\left\|\theta_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\Big{\}}\Big{)}$ (3.12)
$=\sum_{0<|k|\leqslant m}\varepsilon\log\mathbb{E}\Big{(}\exp\Big{\{}\frac{\eta}{\varepsilon}|k|^{2\alpha-4\beta}\langle e_{k},\theta_{\varepsilon,m}(0)\rangle^{2}\Big{\}}\Big{)}$
$=\sum_{0<|k|\leqslant m}\varepsilon\log(1-\eta)^{-1/2}\leqslant\varepsilon\sum_{0<|k|\leqslant m}1\leqslant C(\eta).$
Taking supremum and expectation to (3.11), combining with Hölder’s inequality, (3.10) and (3.12), we obtain
$\mathbb{E}\exp\left\{\frac{\eta}{\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$ (3.13)
$=$ $\mathbb{E}\left[\sup_{t\in[0,T]}\exp\left\{\frac{\eta}{\varepsilon}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\right]$
$\leqslant$ $\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}\mathbb{E}\exp\bigg{[}\left\{\frac{\eta}{\varepsilon}\left\|\theta_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right\}$
$\cdot$ $\sup_{t\in[0,T]}\left(\exp\left\{-\frac{\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right)\bigg{]}$
$\lesssim$ $\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}\left[\mathbb{E}\exp\left\{\frac{2\eta}{\varepsilon}\left\|\theta_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right\}\right]^{1/2}$
$\cdot$ $\left[\mathbb{E}\sup_{t\in[0,T]}\exp\left\{-\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right]^{1/2}$
$\lesssim$ $\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}\left[\mathbb{E}\exp\left\{-\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right]^{1/2}$
By the covariance structure of Brownian motions,
$-\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle$
has the quadratic variation process
$\left\langle\left\langle-\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{\cdot}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\rangle\right\rangle_{t}=\frac{16\eta^{2}}{\varepsilon}\int_{0}^{t}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s.$
Thus $\exp\left\{-\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle-\frac{8\eta^{2}}{\varepsilon}\int_{0}^{t}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$ is a local martingale. Moreover, for any $\eta\in(0,1/8)$,
$M_{2}(t):=\exp\left\{-\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle-\frac{\eta}{\varepsilon}\int_{0}^{t}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$
is a supermartingale. In combination with Hölder’s inequality, we deduce that
$\mathbb{E}\left[\exp\left\{-\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right]$ (3.14)
$\lesssim$ $\left[\mathbb{E}\exp\left\{\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\right]^{1/2}$
$\cdot\left[\mathbb{E}\exp\left\{-\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle-\frac{\eta}{\varepsilon}\int_{0}^{t}\ \left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\right]^{1/2}$
$\lesssim$ $\left[\mathbb{E}\exp\left\{\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\right]^{1/2}.$
Substituting (3.14) into (3.13), it follows that
$\quad\mathbb{E}\exp\left\{\frac{\eta}{2\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$
$\leqslant\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}\left[\mathbb{E}\exp\left\{-\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right]^{1/2}$
$\leqslant\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}.$
In combination with (3.13) and (3.14), we conclude that
$\mathbb{E}\exp\left\{\frac{\eta}{\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{2\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\leqslant\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}.$
Finally, according to the exponential Chebyshev’s inequality,
$\varepsilon\log\mathbb{P}\left(\frac{\eta}{\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{2\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\geqslant R\right)$
$\leqslant$ $-\varepsilon R+\varepsilon\log\mathbb{E}\left(\frac{\eta}{\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{2\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)$
$\leqslant$ $-\varepsilon R+C(\eta)T.$
We complete the proof by sending $R\rightarrow\infty$. ∎
Remark 3.4.
The proof of the exponential estimates for $\theta_{\varepsilon,\delta}$ can be handled in much the same way under the scaling regime (2.7) with initial data satisfying Assumption A. The only difference is that we need to replace the finite-dimensional Itô’s formula by the pathwise $H^{\alpha-2\beta}$-energy equality (3.1). In contrast, we prove the following exponential estimates for $\theta_{\varepsilon,\delta}$ under the scaling regime (2.7). These results hold for $\theta_{\varepsilon,m}$ under the scaling regime (2.8) by the same arguments.
Lemma 3.5.
Let $\varepsilon,\delta>0$. Suppose that $\theta_{\varepsilon,\delta}$ is a stochastic generalized Leray solution of (1.3) in the sense of Definition 3.1, with initial data $\theta_{\varepsilon,\delta}(0)$ satisfying Assumption A. Fix $\sigma\in(0,1/2)$ and $l>3+2\beta-\alpha$. Then there exists $\eta_{0}>0$ such that for every $\eta\in[0,\eta_{0}]$, under the scaling regime (2.7), we have
$\varepsilon\log\mathbb{E}\Big{(}\exp\Big{\{}\frac{\eta}{2\varepsilon}\|\theta_{\varepsilon,\delta}\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}^{2}\Big{\}}\Big{)}\leqslant C(\eta,T),$ (3.15)
and
$\limsup_{R\rightarrow\infty}\sup_{\varepsilon\in(0,1)}\varepsilon\log\mathbb{P}\Big{(}\|\theta_{\varepsilon,\delta}\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}>R\Big{)}=-\infty.$ (3.16)
Proof.
Taking the $W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))$ norm to $\theta_{\varepsilon,\delta}(t)$, it follows from the convexity of the exponential function that
$\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\|\theta_{\varepsilon,\delta}\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big{\}}\lesssim J_{1}+J_{2}+J_{3}+J_{4},$
where
$J_{1}:=$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\|\theta_{\varepsilon,\delta}(0)\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big{\}},$
$J_{2}:=$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\Big{\|}\int^{\cdot}_{0}\Lambda^{2\alpha}\theta_{\varepsilon,\delta}\Big{\|}_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big{\}},$
$J_{3}:=$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\Big{\|}\int^{\cdot}_{0}\nabla\cdot\left(\theta_{\varepsilon,\delta}R^{\perp}\theta_{\varepsilon,\delta}\right)\Big{\|}_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big{\}},$
$J_{4}:=$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon^{1/2}}\|\Lambda^{2\beta}\sqrt{Q_{\delta}}W\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big{\}}.$
Due to (B.1) and the fact that
$\|\theta_{\varepsilon,\delta}(0)\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}\leqslant C(T)\|\theta_{\varepsilon,\delta}(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})},$
we have
$J_{1}\leqslant C(\eta,T).$ (3.17)
Using the Sobolev embedding $W^{1,2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))\subset W^{\sigma,2}([0,T];H^{2\alpha-l}(\mathbb{T}^{2}))$, we deduce that
$J_{2}\lesssim$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\Big{\|}\int^{\cdot}_{0}\Lambda^{\alpha}\theta_{\varepsilon,\delta}\Big{\|}_{W^{1,2}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2}))}\Big{\}}$ (3.18)
$\lesssim$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\int^{T}_{0}\|\Lambda^{\alpha}\theta_{\varepsilon,\delta}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\Big{\}}\leqslant C(\eta,T).$
For the nonlinear term, since $l>3+2\beta-\alpha$, (2.2) and (2.4) imply that for any test function $\phi\in H^{\infty}(\mathbb{T}^{2})$, for any $j\in\mathbb{Z}$ and $\iota>0,$
$\quad|\langle\nabla\cdot\left(\theta_{\varepsilon,\delta}R^{\perp}\theta_{\varepsilon,\delta}\right),\phi\rangle|$ (3.19)
$\lesssim|\langle\theta_{\varepsilon,\delta}R^{\perp}(S_{j}\theta_{\varepsilon,\delta}),\nabla\phi\rangle|+|\langle S_{j}\theta_{\varepsilon,\delta}R^{\perp}(H_{j}\theta_{\varepsilon,\delta}),\nabla\phi\rangle|+|\langle\nabla\cdot\left(H_{j}\theta_{\varepsilon,\delta}R^{\perp}(H_{j}\theta_{\varepsilon,\delta})\right),\phi\rangle|$
$\lesssim\left[\left\|S_{j}\theta_{\varepsilon,\delta}\right\|_{H^{2\beta-\alpha}(\mathbb{T}^{2})}\|\theta_{\varepsilon,\delta}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}+\|\theta_{\varepsilon,\delta}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right]\|\phi\|_{H^{3+2\beta-\alpha+\iota}(\mathbb{T}^{2})}$
$\lesssim\|\theta_{\varepsilon,\delta}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\|\phi\|_{H^{3+2\beta-\alpha+\iota}(\mathbb{T}^{2})}.$
Consequently,
$J_{3}\lesssim\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\int^{T}_{0}\|\theta_{\varepsilon,\delta}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\Big{\}}\leqslant C(\eta,T).$ (3.20)
It remains to estimate the Brownian motion term $J_{4}$. For any $c>0,$ it follows from Young’s inequality that
$J_{4}\lesssim$ $\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon^{1/2}}\|W(\cdot)\|_{W^{\sigma,2}([0,T];H^{\alpha-2\beta-1}(\mathbb{T}^{2}))}\Big{\}}$ (3.21)
$\lesssim$ $\varepsilon\log\mathbb{E}\exp\Big{\{}c\|W(\cdot)\|_{W^{\sigma,2}([0,T];H^{\alpha-2\beta-1}(\mathbb{T}^{2}))}^{2}+\frac{\eta^{2}}{4c\varepsilon}\Big{\}}.$
Since $W$ induces a Gaussian measure on $W^{\sigma,2}([0,T];H^{\alpha-2\beta-1}(\mathbb{T}^{2}))$, with the help of Fernique’s theorem [DPZ14, Remark 2.8], one can choose a sufficiently small $c>0$ such that
$\mathbb{E}\exp\Big{\{}c\|W(\cdot)\|_{W^{\sigma,2}([0,T];H^{\alpha-2\beta-1}(\mathbb{T}^{2}))}^{2}\Big{\}}\leqslant C(T)<\infty.$ (3.22)
Substituting (3.22) into (3.21), we obtain that
$J_{4}\leqslant C(\eta,T).$ (3.23)
Adding (3.17), (3.18), (3.20) and (3.23) together, we complete the proof of (3.15). Finally, (3.16) follows from exponential Chebyshev’s inequality.
Lemma 3.6.
Let $\varepsilon,\delta$ and $\theta_{\varepsilon,\delta}$ satisfy the conditions in Lemma 3.5. Fix $\sigma\in(0,1/2)$ and $l>3+2\beta-\alpha$. Then there exists $\eta_{0}>0$ such that for every $\eta\in[0,\eta_{0}]$, under the scaling regime (2.7), we have
$\varepsilon\log\mathbb{E}\Big{(}\exp\Big{\{}\frac{\eta}{2\varepsilon}\|\theta_{\varepsilon,\delta}\|_{C^{\sigma}([0,T];H^{-l}(\mathbb{T}^{2}))}^{2}\Big{\}}\Big{)}\leqslant C(\eta,T),$
and
$\limsup_{R\rightarrow\infty}\sup_{\varepsilon\in(0,1)}\varepsilon\log\mathbb{P}\Big{(}\|\theta_{\varepsilon,\delta}\|_{C^{\sigma}([0,T];H^{-l}(\mathbb{T}^{2}))}>R\Big{)}=-\infty.$
Proof.
The proof is similar to the proof of Lemma 3.5. We only focus on the estimate of the nonlinear term. For any $\phi\in H^{\infty}(\mathbb{T}^{2})$ and any $0\leqslant s<t\leqslant T$, (3.19) and Hölder’s inequality imply that
$|\langle\int^{t}_{s}\nabla\cdot\left(\theta_{\varepsilon,\delta}R^{\perp}\theta_{\varepsilon,\delta}\right)\,\mathrm{d}r,\phi\rangle|$ $\lesssim\int^{t}_{s}|\langle\nabla\cdot\left(\theta_{\varepsilon,\delta}R^{\perp}\theta_{\varepsilon,\delta}\right),\phi\rangle|\,\mathrm{d}r$
$\lesssim\int^{t}_{s}\|\theta_{\varepsilon,\delta}\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\|\phi\|_{H^{l}(\mathbb{T}^{2})}\,\mathrm{d}r$
$\lesssim(t-s)^{1/2}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,\delta}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\|\phi\|_{H^{l}(\mathbb{T}^{2})}.$
It follows that
$\quad\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\Big{\|}\int^{\cdot}_{0}\nabla\cdot\left(\theta_{\varepsilon,\delta}R^{\perp}\theta_{\varepsilon,\delta}\right)\Big{\|}_{C^{\sigma}([0,T];H^{-l}(\mathbb{T}^{2}))}\Big{\}}$
$\lesssim\varepsilon\log\mathbb{E}\exp\Big{\{}\frac{\eta}{\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,\delta}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\Big{\}}\leqslant C(\eta,T).$
Proposition 3.7 (Exponential tightness).
Let $\varepsilon,\delta$ and $\theta_{\varepsilon,\delta}$ satisfy the conditions in Lemma 3.5. Then under the scaling regime (2.7), there exists a sequence of compact set $\{K_{n}\}_{n\geqslant 1}\subset\mathbb{X}_{\alpha,\beta}$, such that the laws $\{\mu_{\varepsilon,\delta}\}$ of $\{\theta_{\varepsilon,\delta}\}$ satisfy
$\limsup_{n\rightarrow\infty}\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,\delta}(K_{n})=-\infty.$
Proof.
This follows by the same argument as in [GHW24, Corollary 4.2]. In our case, $K_{n}$ can be chosen as
$K_{n}:=\{\theta\in\mathbb{X}_{\alpha,\beta}:\|\theta\|_{L^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))}+\|\theta\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}+\|\theta\|_{C^{\sigma}([0,T];H^{-l}(\mathbb{T}^{2}))}\leqslant n\},$
where $\sigma\in(0,1/2)$ and $l>3+2\beta-\alpha$. Indeed, Lemma 3.3, Lemma 3.5, and Aubin-Lions lemma imply the exponential tightness of $\{\mu_{\varepsilon,\delta}\}$ in $L^{2}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2})).$ The exponential tightness in $L_{w}^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))$ is a direct conclusion of Lemma 3.3. Finally, it follows from Lemma 3.6 that $\{\mu_{\varepsilon,\delta}\}$ is exponentially tight in $C([0,T];H^{\alpha-2\beta}_{w}(\mathbb{T}^{2}))$. ∎
4. Characterization of Rate Function
Let $g\in L^{2}([0,T];H)$. We recall that the skeleton equation for the large deviations is given by (1.1) in our setting, and the rate function is given by
$\mathcal{I}(\theta)=\mathcal{I}_{0}(\theta(0))+\mathcal{I}_{dyna}(\theta),\ \ \theta\in\mathbb{X}_{\alpha,\beta},$ (4.1)
where
$\mathcal{I}_{0}(\phi)=\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2},\quad\phi\in H^{\alpha-2\beta}(\mathbb{T}^{2}),$ (4.2)
and
$\mathcal{I}_{dyna}(\theta)=\frac{1}{2}\inf\Big{\{}\|g\|_{L^{2}([0,T];H)}^{2}:\partial_{t}\theta=\Lambda^{2\alpha}\theta-u_{\theta}\cdot\nabla\theta+\Lambda^{2\beta}g\Big{\}}.$ (4.3)
The equality in (4.3) holds in the following sense.
Definition 4.1 (Weak solution of the skeleton equation).
Let $\theta_{0}\in H^{\alpha-2\beta}(\mathbb{T}^{2})$ and $g\in L^{2}([0,T];H)$. We say that $\theta$ is a weak solution of (1.1) with initial data $\theta_{0}$ if
(i) $\theta\in L^{\infty}([0,T];H^{\alpha-2\beta}(\mathbb{T}^{2}))\cap L^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))\cap C([0,T];H_{w}^{\alpha-2\beta}(\mathbb{T}^{2})).$
(ii) For every $\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))$, for all $t\in[0,T]$,
$\langle\theta(t),\varphi(t)\rangle$ $=\langle\theta_{0},\varphi(0)\rangle+\int^{t}_{0}\langle\theta,\partial_{s}\varphi\rangle\,\mathrm{d}s-\int^{t}_{0}\langle\theta,\Lambda^{2\alpha}\varphi\rangle\,\mathrm{d}s$ (4.4)
$-\frac{1}{2}\int^{t}_{0}\left\langle R_{2}\theta,\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}s+\frac{1}{2}\int_{0}^{t}\left\langle R_{1}\theta,\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}s+\int^{t}_{0}\langle\Lambda^{2\beta}\varphi,g\rangle\,\mathrm{d}s.$
In the sequel, we will provide a variational characterization of $\mathcal{I}$, which is equivalent to the expression (4.1), (4.2) and (4.3). We now define the map $\Lambda_{0}(\cdot,\cdot):C_{b}(H^{\alpha-2\beta}_{w}(\mathbb{T}^{2}))\times H^{\alpha-2\beta}(\mathbb{T}^{2})\rightarrow\mathbb{R}$ by
$\Lambda_{0}(\psi,\phi):=\psi(\phi)-\lambda(\psi),$ (4.5)
where
$\lambda(\psi):=\sup_{\phi\in{H^{\alpha-2\beta}(\mathbb{T}^{2})}}\{\psi(\phi)-\mathcal{I}_{0}(\phi)\}.$ (4.6)
Moreover, $\Lambda_{1}^{T}(\cdot,\cdot):C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))\times\mathbb{X}_{\alpha,\beta}\rightarrow\mathbb{R}$ is defined by
$\Lambda_{1}^{T}(\varphi,$ $\theta):=\langle\theta(T),\varphi(T)\rangle-\langle\theta(0),\varphi(0)\rangle-\int_{0}^{T}\langle\theta,\partial_{t}\varphi\rangle\,\mathrm{d}t+\int_{0}^{T}\langle\theta,\Lambda^{2\alpha}\varphi\rangle\,\mathrm{d}t$
$+\frac{1}{2}\int_{0}^{T}\left\langle R_{2}\theta,\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}t-\frac{1}{2}\int_{s}^{t}\left\langle R_{1}\theta,\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}t-\frac{1}{2}\|\Lambda^{2\beta}\varphi\|_{L^{2}([0,T];H)}^{2}.$
Proposition 4.2.
Let the rate functions $\mathcal{I}_{0},\mathcal{I}_{dyna},\mathcal{I}$ be defined by (4.2), (4.3) and (4.1), respectively. Then for any $\theta\in\mathbb{X}_{\alpha,\beta}$ with $\mathcal{I}(\theta)<\infty$, we have
$\mathcal{I}(\theta)=\sup_{\begin{subarray}{c}\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))\\ \psi\in C_{b}(H^{\alpha-2\beta}_{w}(\mathbb{T}^{2}))\end{subarray}}\Big{\{}\Lambda_{0}(\psi,\theta(0))+\Lambda_{1}(\varphi,\theta)\Big{\}}.$ (4.7)
Proof.
It is sufficient to prove that for any $\theta\in\mathbb{X}_{\alpha,\beta}$ with $\mathcal{I}(\theta)<\infty$, we have
$\mathcal{I}_{0}(\theta(0))=\sup_{\psi\in C_{b}(H^{\alpha-2\beta}_{w}(\mathbb{T}^{2}))}\Lambda_{0}(\psi,\theta(0)),$ (4.8)
and
$\mathcal{I}_{dyna}(\theta)=\sup_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta).$ (4.9)
According to Proposition B.1, $\mu_{\varepsilon,\delta}^{0}=\mathcal{G}(0,\varepsilon Q_{\delta}/2)$ satisfies the large deviations on $H^{\alpha-2\beta}_{w}(\mathbb{T}^{2})$ with rate function $\mathcal{I}_{0}$ under Assumption A and the scaling regime (2.7). Thus (4.8) follows from Proposition B.1 and Bryc’s lemma (see, for example, [DZ10, Theorem 4.4.2]). Next we prove (4.9). Let $F^{T}(\cdot,\theta):C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))\rightarrow\mathbb{R}$ be the linear part of $\Lambda_{1}^{T}(\cdot,\theta):$
$F^{T}(\varphi,\theta):=$ $\langle\theta(T),\varphi(T)\rangle-\langle\theta(0),\varphi(0)\rangle-\int_{0}^{T}\langle\theta,\partial_{t}\varphi\rangle\,\mathrm{d}t+\int_{0}^{T}\langle\theta,\Lambda^{2\alpha}\varphi\rangle\,\mathrm{d}t$ (4.10)
$+\frac{1}{2}\int_{0}^{T}\left\langle R_{2}\theta,\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}t-\frac{1}{2}\int_{0}^{T}\left\langle R_{1}\theta,\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}t.$
For any $\theta$ satisfying $I_{dyna}(\theta)<\infty$, by the definition of $\mathcal{I}_{dyna}(\theta)$, there exists a function $g\in L^{2}([0,T];H)$ such that, for any $\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))$,
$F^{T}(\varphi,\theta)=\int_{0}^{T}\langle g,\Lambda^{2\beta}\varphi\rangle\,\mathrm{d}t.$
Using the definition of $\Lambda_{1}^{T}(\cdot,\theta)$ and Cauchy-Schwarz inequality, we have
$\Lambda_{1}^{T}(\varphi,\theta)=\int^{T}_{0}\langle g,\Lambda^{2\beta}\varphi\rangle\,\mathrm{d}t-\frac{1}{2}\|\Lambda^{2\beta}\varphi\|_{L^{2}([0,T];H)}^{2}\leqslant\frac{1}{2}\|g\|_{L^{2}([0,T];H)}^{2}.$
Taking the supremum over $\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))$, it follows that
$\sup_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta)\leqslant\frac{1}{2}\|g\|_{L^{2}([0,T];H)}^{2}.$ (4.11)
Taking the infimum of the right-hand side of (4.11) over all $g\in L^{2}([0,T];H)$ such that $\theta$ is a weak solution of (1.1) with $g$, we conclude that
$\sup_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta)\leqslant\mathcal{I}_{dyna}(\theta).$ (4.12)
In the sequel, we prove the reverse inequality of (4.12). For the case $\sup\limits_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta)=\infty$, it is clear that
$\mathcal{I}_{dyna}(\theta)\leqslant\sup\limits_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta).$
Therefore it is sufficient to discuss the case $\sup\limits_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta)<\infty$. By the definition of $\Lambda_{1}^{T}$,
$\sup_{\begin{subarray}{c}\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2})),\\ \|\varphi\|_{L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2}))}\leqslant 1\end{subarray}}|F^{T}(\varphi,\theta)|\leqslant$ $\sup\limits_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta)+\frac{1}{2}<+\infty.$
Since $C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))$ is dense in $L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2}))$, $F^{T}(\cdot,\theta)$ can be extended to a continuous linear functional from $L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2}))$ to $\mathbb{R}$. Using Riesz’s representation theorem, one can see that there exists $\Psi^{\theta}\in L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2}))$ such that
$F^{T}(\varphi,\theta)=\langle\Psi^{\theta},\varphi\rangle_{L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2}))}.$ (4.13)
Let $g=g^{\theta}:=\Lambda^{2\beta}\Psi^{\theta}\in L^{2}([0,T];H)$, (4.13) implies that $\theta$ is a weak solution of (1.1) with respect to the control $g$ in the sense of Definition 4.1. Moreover,
$\|g^{\theta}\|_{L^{2}([0,T];H)}=\|\Psi\|_{L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2}))}=\|F^{T}(\cdot,\theta)\|_{(L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2})))^{\prime}}.$
Here $(L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2})))^{\prime}$ is the dual space of the Hilbert space $L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2})).$ From the definition of $F^{T}$ and $\Lambda_{1}^{T}$, we have
$\Lambda_{1}^{T}(\varphi,\theta)=F^{T}(\varphi,\theta)-\frac{1}{2}\|\Lambda^{2\beta}\varphi\|_{L^{2}([0,T];H)}^{2}.$
Thanks to the fact that $C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))$ is dense in $L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2}))$, and the calculation of the extremum of a quadratic functional on a Hilbert space, it follows that
$\sup_{\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))}\Lambda_{1}^{T}(\varphi,\theta)=$ $\sup_{\varphi\in L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2}))}\Big{(}F^{T}(\varphi,\theta)-\frac{1}{2}\|\Lambda^{2\beta}\varphi\|_{L^{2}([0,T];H)}^{2}\Big{)}$
$=$ $\frac{1}{2}\|F^{T}(\cdot,\theta)\|_{(L^{2}([0,T];H^{2\beta}(\mathbb{T}^{2})))^{\prime}}^{2}=\frac{1}{2}\|g^{\theta}\|_{L^{2}([0,T];H)}^{2}\geqslant\mathcal{I}_{dyna}(\theta).$
This completes the proof. ∎
Remark 4.3.
It can be seen from the above proof that, for any $\theta\in\mathbb{X}_{\alpha,\beta}$ with $\mathcal{I}(\theta)<\infty$, $g^{\theta}$ is the unique element in $L^{2}([0,T];H)$ such that $\theta$ is a weak solution of (1.1).
Lemma 4.4.
Let $\mathcal{I}$ be the rate function defined by (4.1). Then $\mathcal{I}$ is lower semi-continuous with respect to the topology of $\mathbb{X}_{\alpha,\beta}$.
Proof.
The lower semi-continuity of the $H^{\alpha-2\beta}(\mathbb{T}^{2})$-norm with respect to the weak topology of $H^{\alpha-2\beta}(\mathbb{T}^{2})$ implies that the map $\theta\mapsto\mathcal{I}_{0}(\theta(0))$ is lower semi-continuous on $\mathbb{X}_{\alpha,\beta}$. We are left with the task of proving the lower semi-continuity of $\mathcal{I}_{dyna}$ Suppose that $\theta\in\mathbb{X}_{\alpha,\beta}$ and $\{\theta_{n}\}_{n\geqslant 1}\subset\mathbb{X}_{\alpha,\beta}$ satisfy $\theta_{n}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\beta}$ as $n\rightarrow\infty$. With the help of Proposition 4.2, it is sufficient to show that for arbitrary fixed $\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))$, the map $\Lambda_{1}^{T}(\varphi,\cdot)$ is continuous with respect to the topology of $\mathbb{X}_{\alpha,\beta}$:
$\lim_{n\rightarrow\infty}\Lambda_{1}^{T}(\varphi,\theta_{n})=\Lambda_{1}^{T}(\varphi,\theta).$ (4.14)
The definition of the topology of $\mathbb{X}_{\alpha,\beta}$ implies the convergence of $\langle\theta_{n}(T),\varphi(T)\rangle$, $\langle\theta_{n}(0),\theta(0)\rangle$ and $\int_{0}^{T}\langle\theta_{n},\partial_{t}\varphi\rangle\,\mathrm{d}t$ directly. The convergence of the dissipation term $\int_{0}^{T}\langle\theta_{n},\Lambda^{2\alpha}\varphi\rangle\,\mathrm{d}t$ follows from the fact that $\theta_{n}\rightarrow\theta$ weakly in $L^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2}))$.
It remains to show the convergence of the nonlinear term. For any $j\in\mathbb{Z}$, according to (2.2) and (2.3), it holds that
$\Big{|}\int_{0}^{T}\left(\left\langle R_{2}\theta_{n},\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta_{n}\right\rangle-\left\langle R_{1}\theta_{n},\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta_{n}\right\rangle\right)\mathrm{d}t$
$-\int_{0}^{T}\left(\left\langle R_{2}\theta,\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta\right\rangle-\left\langle R_{1}\theta,\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta\right\rangle\right)\mathrm{d}t$ $\Big{|}\leqslant I^{n}_{1j}+I^{n}_{2j}+I^{n}_{3j}+I_{4j},$
where
$I^{n}_{1j}$ $:=\left|\int_{0}^{T}\langle\nabla\cdot\left(\theta_{n}R^{\perp}\left(S_{j}\theta_{n}\right)\right),\varphi\rangle\,\mathrm{d}t-\int_{0}^{T}\langle\nabla\cdot\left(\theta R^{\perp}\left(S_{j}\theta\right)\right),\varphi\rangle\,\mathrm{d}t\right|,$
$I^{n}_{2j}$ $:=\left|\int_{0}^{T}\langle\nabla\cdot\left(S_{j}\theta_{n}R^{\perp}\left(H_{j}\theta_{n}\right)\right),\varphi\rangle\,\mathrm{d}t-\int_{0}^{T}\langle\nabla\cdot\left(S_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right),\varphi\rangle\,\mathrm{d}t\right|,$
$I^{n}_{3j}$ $:=\left|\int_{0}^{T}\langle\nabla\cdot\left(H_{j}\theta_{n}R^{\perp}\left(H_{j}\theta_{n}\right)\right),\varphi\rangle\,\mathrm{d}t\right|,$
and
$I_{4j}:=\left|\int_{0}^{T}\langle\nabla\cdot\left(H_{j}\theta R^{\perp}\left(H_{j}\theta\right)\right),\varphi\rangle\,\mathrm{d}t\right|.$
The convergence of $\theta_{n}$ in $\mathbb{X}_{\alpha,\beta}$ implies that the $L^{2}\left([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2})\right)$-norm of $\{\theta_{n}\}$ is bounded:
$\int_{0}^{T}\left\|\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}t\leqslant C.$ (4.15)
Using (2.4) for $r=2\beta-2\alpha$, (4.15) and the fact that $\left\|H_{j}\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}\leqslant\left\|\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})},$ we have
$I^{n}_{3j}$ $\leqslant C(\varphi)\int_{0}^{T}\left\|H_{j}\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}\left\|H_{j}\theta_{n}\right\|_{H^{2\beta-2\alpha-1}(\mathbb{T}^{2})}\,\mathrm{d}t$ (4.16)
$\lesssim C(\varphi)2^{-j(4\alpha+1-4\beta)}\int_{0}^{T}\left\|H_{j}\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}t$
$\lesssim C(\varphi)2^{-j(4\alpha+1-4\beta)}\int_{0}^{T}\left\|\theta_{n}\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}t$
$\lesssim C(\varphi)2^{-j(4\alpha+1-4\beta)}.$
Applying the same argument as in $I^{n}_{3j}$, we can see that,
$\lim\limits_{j\rightarrow\infty}I_{4j}=0.$ (4.17)
According to (4.16) and (4.17), for any $\eta>0,$ we can choose a sufficiently large positive integer $j_{0},$ such that $I^{n}_{3j_{0}}+I_{4j_{0}}\leqslant\eta$ for all $n\geqslant 1.$ By [Mar08a, Lemma 9.3], $S_{j_{0}}\theta_{n}\rightarrow S_{j_{0}}\theta$ and $R^{\perp}(S_{j_{0}}\theta_{n})\rightarrow R^{\perp}(S_{j_{0}}\theta)$ strongly in $L^{2}([0,T];H^{1/2}(\mathbb{T}^{2}))$, which implies
$\lim\limits_{n\rightarrow\infty}I^{n}_{1j_{0}}=\lim\limits_{n\rightarrow\infty}I^{n}_{2j_{0}}=0$ (4.18)
since $\theta_{n}\rightarrow\theta$ weakly in $L^{2}([0,T];H^{2\alpha-2\beta}(\mathbb{T}^{2})).$ Hence
$\limsup_{n\rightarrow\infty}\left(I^{n}_{1j_{0}}+I^{n}_{2j_{0}}+I^{n}_{3j_{0}}+I_{4j_{0}}\right)\leqslant\eta.$
By the arbitrariness of $\eta,$ we conclude that
$\lim_{n\rightarrow\infty}$ $\int_{0}^{T}\left\langle R_{2}\theta_{n},\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta_{n}\right\rangle-\left\langle R_{1}\theta_{n},\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta_{n}\right\rangle\mathrm{d}t$
$=\int_{0}^{T}\left\langle R_{2}\theta,\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta\right\rangle-\left\langle R_{1}\theta,\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta\right\rangle\mathrm{d}t.$
5. Upper Bound for Large Derivations
In this section, we proceed with the proof of the large deviations for (1.3). Thanks to the exponential tightness (Proposition 3.7), it suffices to prove the weak large deviations in the sense that the upper bound holds for all compact sets in $\mathbb{X}_{\alpha,\beta}$.
For every $\psi\in C_{b}(H_{w}^{\alpha-2\beta}(\mathbb{T}^{2}))$ and $\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))$, we define $\mathcal{M}^{\psi,\varphi}:[0,T]\times\mathbb{X}_{\alpha,\beta}\rightarrow\mathbb{R}$ by
$\mathcal{M}^{\psi,\varphi}(t,\theta):=F^{t}(\varphi,\theta)+\Lambda_{0}(\psi,\theta(0)),$
where $F^{t}(\varphi,\theta)$ is defined by (4.10) with $T$ replaced by $t$. Let $Q^{\psi,\varphi}:[0,T]\times\mathbb{X}_{\alpha,\beta}\rightarrow\mathbb{R}$ be defined by
$Q^{\psi,\varphi}(t,\theta):=\exp\Big{\{}\varepsilon^{-1}\big{(}\mathcal{M}^{\psi,\varphi}(t,\theta)-\frac{1}{2}\int^{t}_{0}\|\varphi\|_{H^{2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\big{)}\Big{\}}.$
Proposition 5.1 (Upper bound).
For every $\varepsilon,\delta>0$ and $m\in\mathbb{N}_{+}$, let $\theta_{\varepsilon,\delta}$ be a stochastic generalized Leray solution of (1.3) in the sense of Definition 3.1 with initial data $\theta_{\varepsilon,\delta}(0)$ satisfying Assumption A, and let $\theta_{\varepsilon,m}$ be a solution of (3.7) with initial data $\theta_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$.
(i) Let $\mu_{\varepsilon}=\mu_{\varepsilon,\delta}$ be the laws of $\theta_{\varepsilon,\delta}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.7) holds for $(\varepsilon,\delta)$, then for any closed set $F\subset\mathbb{X}_{\alpha,\beta}$,
$\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon}(F)\leqslant-\inf_{\theta\in F}\mathcal{I}(\theta).$ (5.1)
(ii) Let $\mu_{\varepsilon}=\mu_{\varepsilon,m}$ be the laws of $\theta_{\varepsilon,m}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.8) holds for $(\varepsilon,m)$, then for any closed set $F\subset\mathbb{X}_{\alpha,\beta}$, the upper bound (5.1) holds as well.
Proof.
We only prove (i), and (ii) could be handled in a similar way as (i). Due to Proposition 3.7, it suffices to prove (5.1) for every compact $F$ in $\mathbb{X}_{\alpha,\beta}$.
By Assumption A, Proposition B.1 and Varadhan’s integration lemma (see, for example, [DZ10, Theorem 4.3.1]),
$\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\int_{\mathbb{X}_{\alpha,\beta}}Q^{\psi,\varphi}(0,\theta)\mu_{\varepsilon}(\mathrm{d}\theta)$ $=\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\int_{\mathbb{X}_{\alpha,\beta}}\exp\{\varepsilon^{-1}\Lambda_{0}(\psi,\theta(0))\}\mu_{\varepsilon}(\mathrm{d}\theta)$ (5.2)
$=\limsup_{\varepsilon\rightarrow 0}\Big{(}\varepsilon\log\mathbb{E}\exp\{\varepsilon^{-1}\psi(\theta_{\varepsilon}(0))\}-\lambda(\psi)\Big{)}=0.$
Since $\theta_{\varepsilon,\delta}$ is a solution of (1.3) in the sense of Definition 3.1 with initial data $\theta_{\varepsilon,\delta}(0)$, the stochastic process $\mathcal{M}^{\psi,\varphi}(t,\theta_{\varepsilon,\delta})$ is a martingale. Due to the contraction property of $Q_{\delta}$ on $H^{2\beta}(\mathbb{T}^{2})$, the quadratic variation of $\mathcal{M}^{\psi,\varphi}(\cdot,\theta_{\varepsilon,\delta})$ satisfies
$\langle\langle\mathcal{M}^{\psi,\varphi}(\cdot,\theta_{\varepsilon,\delta})\rangle\rangle_{t}=\varepsilon\int_{0}^{t}\|\sqrt{Q_{\delta}}\Lambda^{2\beta}\varphi\|_{H}^{2}\,\mathrm{d}s\leqslant\varepsilon\int_{0}^{t}\|\varphi\|_{H^{2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s.$
It follows that $Q^{\psi,\varphi}(t,\theta_{\varepsilon,\delta})$ is a supermartingale. For every compact set $F\subset\mathbb{X}_{\alpha,\beta}$, we have
$\mu_{\varepsilon}(F)\leqslant$ $\sup_{\theta\in F}(Q^{\psi,\varphi}(T,\theta))^{-1}\int_{F}Q^{\psi,\varphi}(T,\theta)\mu_{\varepsilon}(\mathrm{d}\theta)$ (5.3)
$\leqslant$ $\sup_{\theta\in F}(Q^{\psi,\varphi}(T,\theta))^{-1}\int_{\mathbb{X}_{\alpha,\beta}}Q^{\psi,\varphi}(T,\theta)\mu_{\varepsilon}(\mathrm{d}\theta)$
$\leqslant$ $\sup_{\theta\in F}(Q^{\psi,\varphi}(T,\theta))^{-1}\int_{\mathbb{X}_{\alpha,\beta}}Q^{\psi,\varphi}(0,\theta)\mu_{\varepsilon}(\mathrm{d}\theta)$
$\leqslant$ $\exp\Big{\{}-\varepsilon^{-1}\inf_{\theta\in F}\Big{(}\mathcal{M}^{\psi,\varphi}(T,\theta)-\frac{1}{2}\int^{T}_{0}\|\varphi\|_{H^{2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\Big{)}\Big{\}}\int_{\mathbb{X}_{\alpha,\beta}}Q^{\psi,\varphi}(0,\theta)\mu_{\varepsilon}(\mathrm{d}\theta).$
In (5.3), we used the non-negativity of $Q^{\psi,\varphi}$ in the first and second inequalities, and properties of supermartingales have been employed in the third inequality. The definition of $Q^{\psi,\varphi}$ has been used to derive the last inequality.
Taking the logarithmic function in (5.3) yields
$\varepsilon\log\mu_{\varepsilon}(F)\leqslant$ $-\inf_{\theta\in F}\Big{\{}\Lambda^{T}_{1}(\varphi,\theta)+\Lambda_{0}(\psi,\theta(0))\Big{\}}+\varepsilon\log\int_{\mathbb{X}_{\alpha,\beta}}Q^{\psi,\varphi}(0,\theta)\mu_{\varepsilon}(\mathrm{d}\theta).$
Letting $\varepsilon\rightarrow 0$ and using (5.2), we conclude that
$\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon}(F)\leqslant$ $\inf_{\begin{subarray}{c}\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))\\ \psi\in C_{b}(H_{w}^{\alpha-2\beta}(\mathbb{T}^{2}))\end{subarray}}\Big{(}-\inf_{\theta\in F}\Big{\{}\Lambda_{1}^{T}(\varphi,\theta)+\Lambda_{0}(\psi,\theta(0))\Big{\}}\Big{)}$
$=$ $-\sup_{\begin{subarray}{c}\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))\\ \psi\in C_{b}(H_{w}^{\alpha-2\beta}(\mathbb{T}^{2}))\end{subarray}}\inf_{\theta\in F}\Big{\{}\Lambda^{T}_{1}(\varphi,\theta)+\Lambda_{0}(\psi,\theta(0))\Big{\}}.$
Since $F$ is compact in $\mathbb{X}_{\alpha,\beta}$, by [KL99, Appendix 2, Lemma 3.2], one can swap the above infimum and supremum. Finally, Proposition 4.2 implies that
$\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon}(F)\leqslant-\inf_{\theta\in F}\mathcal{I}(\theta).$
6. Lower Bound for Large Derivations
In this section, we will prove a restricted large deviations lower bound using the entropy method as described in [Mar10] (see Lemma 2.5). Due to the lack of well-posedness of the skeleton equation (1.1), it is challenging to match the upper and lower bounds in $\mathbb{X}_{\alpha,\beta}$. Instead, we will match the upper and lower bounds within the following weak-strong uniqueness class.
Definition 6.1 (Weak-Strong Uniqueness Class).
We say $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ is a weak-strong uniqueness class of (1.1) for every $g\in L^{2}([0,T];H)$ if, for arbitrary two weak solutions $\theta_{1},\theta_{2}$ of the skeleton equation (1.1) in the sense of Definition 4.1 with the same initial data $\theta_{1}(0)=\theta_{2}(0)$ and the same control $g$, we have $\theta_{1}=\theta_{2}$ in $\mathbb{X}_{\alpha,\beta}$ as long as $\theta_{1}\in\mathcal{C}_{0}$ and $\theta_{2}$ satisfies the $H^{\alpha-2\beta}$-energy inequality: for every $t\in[0,T],$
$\frac{1}{2}\|\theta_{2}(t)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\|\theta_{2}(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\left\|\theta_{2}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\langle\Lambda^{2\alpha-2\beta}\theta_{2},g\rangle\,\mathrm{d}s.$ (6.1)
Definition 6.2 ($\mathcal{I}$-closure).
Suppose that $\mathcal{A}$ is a subset of $\mathbb{X}_{\alpha,\beta}$ and $\mathcal{I}:\mathbb{X}_{\alpha,\beta}\rightarrow[0,+\infty]$ is lower semi-continuous. We define the $\mathcal{I}$-closure of $\mathcal{A}$ as
$\overline{\mathcal{A}}^{\mathcal{I}}:=\big{\{}\theta\in\mathbb{X}_{\alpha,\beta}:\text{there exists a sequence }\{\theta^{(n)}\}\subset\mathcal{A},\,\text{such that }\theta^{(n)}\rightarrow\theta\text{ in }\mathbb{X}_{\alpha,\beta},\,\mathcal{I}\big{(}\theta^{(n)}\big{)}\rightarrow\mathcal{I}(\theta)\big{\}}.$
Let $\mathcal{C}_{0}$ be a weak-strong uniqueness class of (1.1) and let $\mathcal{C}$ be its $\mathcal{I}$-closure. We will take $E=\mathbb{X}_{\alpha,\beta}$ and $I=\mathcal{I}|_{\mathcal{C}_{0}}$ in Lemma 2.5 to obtain the lower bound of large deviations. Moreover, when utilizing Lemma 2.5, we focus only on the case $\theta\in\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ with $\mathcal{I}(\theta)<\infty$. Otherwise, the condition (ii) in Lemma 2.5 holds obviously, and one can take $\pi^{\varepsilon,\theta}=\delta_{\theta}$ to verify the condition (i).
Proposition 6.3 (Restricted lower bound).
For every $\varepsilon,\delta>0$ and $m\in\mathbb{N}_{+}$, let $\theta_{\varepsilon,\delta}$ be a stochastic generalized Leray solution of (1.3) in the sense of Definition 3.1 with initial data $\theta_{\varepsilon,\delta}(0)$ satisfying Assumption A, and let $\theta_{\varepsilon,m}$ be a solution of (3.7) with initial data $\theta_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$. Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ be a weak-strong uniqueness class of (1.1) and let $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$.
(i) Let $\mu_{\varepsilon}=\mu_{\varepsilon,\delta}$ be the law of $\theta_{\varepsilon,\delta}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.7) holds for $(\varepsilon,\delta)$, then for any open set $G\subset\mathbb{X}_{\alpha,\beta}$,
$\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon}(G)\geqslant-\inf_{\theta\in G\cap\mathcal{C}}\mathcal{I}(\theta).$ (6.2)
(ii) Let $\mu_{\varepsilon}=\mu_{\varepsilon,m}$ be the law of $\theta_{\varepsilon,m}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.8) holds for $(\varepsilon,m)$, then for any open set $G\subset\mathbb{X}_{\alpha,\beta}$, the restricted lower bound (6.2) holds as well.
Lemma 6.4.
For every $\theta\in\mathcal{C}_{0}$ with $\mathcal{I}(\theta)<\infty$, let $g\in L^{2}([0,T];H)$ be the control such that $\theta$ solves (1.1) weakly. For every $\varepsilon>0$ and $m\in\mathbb{N}_{+}$ satisfying the scaling regime (2.8), let $\theta_{\varepsilon,m}$ and $\mu_{\varepsilon,m}$ be as in Proposition 6.3. Then there exists a family of probability measure $\{\pi_{\varepsilon,m,\theta}\}$ on $\mathbb{X}_{\alpha,\beta}$ that satisfies the entropy inequality
$\limsup_{\varepsilon\rightarrow 0}\operatorname{Ent}\left(\pi_{\varepsilon,m,\theta}\mid\mu_{\varepsilon,m}\right)\leqslant\mathcal{I}(\theta).$ (6.3)
Proof.
Let $\left(\Omega,\mathcal{F},\{\mathcal{F}(t)\}_{t\in[0,T]},\overline{\mathbb{P}}\right)$ be a stochastic basis and $\overline{W}$ be a cylindrical Wiener process on $H^{\alpha-2\beta}(\mathbb{T}^{2})$ such that $\bar{\theta}_{\varepsilon,m}$ is a solution of (3.7) with initial data $\bar{\theta}_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$ and $\bar{\theta}_{\varepsilon,m}(0)$ is independent of $\overline{W}$. For any two measures $\mu,\nu$ on $\Omega$, we denote by $\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$ the Radon-Nikodym derivative of $\mu$ with respect to $\nu.$ Now we define the $H$-valued random variable $Y_{0}^{\varepsilon,m}$ by
$Y_{0}^{\varepsilon,m}:=\frac{\mathrm{d}\mathcal{G}(P_{m}\theta(0),\varepsilon P_{m}/2)}{\mathrm{d}\mathcal{G}(0,\varepsilon P_{m}/2)}\left(\bar{\theta}_{\varepsilon,m}(0)\right),$
and define the stochastic process $Z^{\varepsilon,m}$ by
$Z^{\varepsilon,m}(t):=\exp\left(-\frac{1}{\sqrt{\varepsilon}}\int_{0}^{t}\left\langle P_{m}g(s),\,\mathrm{d}\overline{W}(s)\right\rangle-\frac{1}{2\varepsilon}\int_{0}^{t}\left\|P_{m}g(s)\right\|_{H}^{2}\,\mathrm{d}s\right)Y_{0}^{\varepsilon,m},\quad t\in[0,T].$ (6.4)
Let $\mathbb{P}$ be the probability measure on $\left(\Omega,\mathcal{F}\right)$ such that $Z^{\varepsilon,m}(T)=\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\overline{\mathbb{P}}}.$ By Cameron-Martin theorem, $W(t):=\overline{W}(t)+\varepsilon^{-1/2}\int_{0}^{t}P_{m}g(s)\,\mathrm{d}s$ is a cylindrical Wiener process on $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ and $\bar{\theta}_{\varepsilon,m}$ is a solution of
$\mathrm{d}\bar{\theta}_{\varepsilon,m}(t)=-\Lambda^{2\alpha}\bar{\theta}_{\varepsilon,m}(t)\,\mathrm{d}t-P_{m}\left(u_{\bar{\theta}_{\varepsilon,m}}(t)\cdot\nabla\bar{\theta}_{\varepsilon,m}(t)\right)\,\mathrm{d}t+\sqrt{\varepsilon}\Lambda^{2\beta}\,\mathrm{d}W^{m}(t)+P_{m}\Lambda^{2\beta}g\,\mathrm{d}t$ (6.5)
with the initial data $\bar{\theta}_{\varepsilon,m}(0)\sim\mathcal{G}(P_{m}\theta(0),\varepsilon P_{m}/2).$
Let $\pi_{\varepsilon,m,\theta}$ be the law of $\bar{\theta}_{\varepsilon,m}$ on $\left(\Omega,\mathcal{F},\mathbb{P}\right)$. According to (6.4), the definition of relative entropy, the expression of $\mathcal{I}_{0}$, and the proof for the lower bound of Proposition B.1,
$\varepsilon\operatorname{Ent}\left(\pi_{\varepsilon,m,\theta}|\mu_{\varepsilon,m}\right)\leqslant\varepsilon\operatorname{Ent}(\mathbb{P}|\overline{\mathbb{P}})$ $=\varepsilon\mathbb{E}\left[\log Z^{\varepsilon,m}(T)\right]$
$=\frac{1}{2}\int_{0}^{T}\|P_{m}g(s)\|_{H}^{2}\,\mathrm{d}s+\varepsilon\mathbb{E}\left[\log Y_{0}^{\varepsilon,m}\right]$
$\leqslant\frac{1}{2}\int_{0}^{T}\|g(s)\|_{H}^{2}\,\mathrm{d}s+\mathcal{I}_{0}(\theta(0))=\mathcal{I}(\theta).$
Lemma 6.5.
For every $\varepsilon>0$ and $m\in\mathbb{N}_{+}$ satisfying the scaling regime (2.8), let $\theta,g,\pi^{\varepsilon,m,\theta}$ and $\mu_{\varepsilon,m}$ be as in Lemma 6.4. Then $\pi^{\varepsilon,m,\theta}$ converges to $\delta_{\theta}$ in law on $\mathbb{X}_{\alpha,\beta}$ as $\varepsilon\rightarrow 0$.
Proof.
The proof will be divided into three steps. First, we show that $\{\pi^{\varepsilon,m,\theta}\}$ is tight in $\mathbb{X}_{\alpha,\beta}$ and that any limit point $\bar{\theta}$ is a weak solution of (1.1) in the sense of Definition 4.1. Next, we prove the $H^{\alpha-2\beta}$-energy inequality (6.1) for $\bar{\theta}$. Finally, we conclude that $\bar{\theta}=\theta$ in $\mathbb{X}_{\alpha,\beta}$ by using the weak-strong uniqueness.
Step 1. $\bar{\theta}$ is a weak solution of the skeleton equation.
Tightness of $\{\bar{\theta}_{\varepsilon,m}\}$ in $\mathbb{X}_{\alpha,\beta}$ follows by the same method as in the proof of Proposition 3.7, thus we omit the proof. According to Jakubowski-Skorokhod representation theorem [Jak97], for any sequence $\varepsilon_{k}\rightarrow 0$, there exists a subsequence $(\varepsilon_{k},m_{k})$ (still denoted by $(\varepsilon,m)$), a new stochastic basis (still denoted by $(\Omega,\mathcal{F},\left\{\mathcal{F}_{t}\right\}_{t\in[0,T]},\mathbb{P})$), a cylindrical Wiener process (still denoted by $W$), a sequence of $\mathbb{X}_{\alpha,\beta}$-valued random elements with the same laws as $\{\bar{\theta}_{\varepsilon,m}\}$ on the new probability space (still denoted by $\{\bar{\theta}_{\varepsilon,m}\}$) and a limit point $\bar{\theta}$, such that $\mathbb{P}$-almost surely,
$\bar{\theta}_{\varepsilon,m}\rightarrow\bar{\theta}\,\text{ in }\mathbb{X}_{\alpha,\beta}.$
We claim that $\bar{\theta}$ is a weak solution of the (1.1). For any $\varphi\in C^{\infty}\left([0,T];H^{\infty}(\mathbb{T}^{2})\right)$ and $t\in[0,T]$, we will pass to the limits in all terms of
$\left\langle\bar{\theta}_{\varepsilon,m}(t),\varphi(t)\right\rangle$ $+\int_{0}^{t}\left\langle\bar{\theta}_{\varepsilon,m},\Lambda^{2\alpha}\varphi\right\rangle\,\mathrm{d}s=\left\langle\bar{\theta}_{\varepsilon,m}(0),\varphi(0)\right\rangle+\int_{0}^{t}\langle\bar{\theta}_{\varepsilon,m},\partial_{s}\varphi\rangle\,\mathrm{d}s$ (6.6)
$-\frac{1}{2}\int^{t}_{0}\left\langle R_{2}\bar{\theta}_{\varepsilon,m},\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\bar{\theta}_{\varepsilon,m}\right\rangle\,\mathrm{d}s+\frac{1}{2}\int_{0}^{t}\left\langle R_{1}\bar{\theta}_{\varepsilon,m},\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\bar{\theta}_{\varepsilon,m}\right\rangle\,\mathrm{d}s$
$+\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\beta}\varphi,P_{m}\,\mathrm{d}W(s)\rangle+\int_{0}^{t}\left\langle P_{m}g,\Lambda^{2\beta}\varphi\right\rangle\,\mathrm{d}s.$
An analysis similar to that in the proof of Lemma 4.4 shows that, $\mathbb{P}$-almost surely, for all $t\in[0,T],$ as $\varepsilon\rightarrow 0,$
$\left\langle\bar{\theta}_{\varepsilon,m}(t),\varphi(t)\right\rangle$ $\rightarrow\left\langle\bar{\theta}(t),\varphi(t)\right\rangle,\quad\int_{0}^{t}\left\langle\bar{\theta}_{\varepsilon,m},\Lambda^{2\alpha}\varphi\right\rangle\,\mathrm{d}s\rightarrow\int_{0}^{t}\left\langle\bar{\theta},\Lambda^{2\alpha}\varphi\right\rangle\,\mathrm{d}s,$
$\left\langle\bar{\theta}_{\varepsilon,m}(0),\varphi(0)\right\rangle$ $\rightarrow\left\langle\bar{\theta}(0),\varphi(0)\right\rangle,\quad\int_{0}^{t}\langle\bar{\theta}_{\varepsilon,m},\partial_{s}\varphi\rangle\,\mathrm{d}s\rightarrow\int_{0}^{t}\langle\bar{\theta},\partial_{s}\varphi\rangle\mathrm{d}s,$
and
$\int^{t}_{0}\left\langle R_{2}\bar{\theta}_{\varepsilon,m},\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\bar{\theta}_{\varepsilon,m}\right\rangle\,$ $\mathrm{d}s-\int_{0}^{t}\left\langle R_{1}\bar{\theta}_{\varepsilon,m},\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\bar{\theta}_{\varepsilon,m}\right\rangle\,\mathrm{d}s\rightarrow$
$-\int^{t}_{0}\left\langle R_{2}\bar{\theta},\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\bar{\theta}\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle R_{1}\bar{\theta},\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\bar{\theta}\right\rangle\,\mathrm{d}s.$
It remains to show the convergence of the last two terms in (6.6). According to Doob’s $L^{2}$ inequality and Itô isometry, it holds that
$\mathbb{E}\left(\sup_{0\leqslant t\leqslant T}\left|\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\beta}\varphi(s),P_{m}\,\mathrm{d}W(s)\rangle\right|^{2}\right)\lesssim\varepsilon\left\|\Lambda^{2\beta}\varphi\right\|_{L^{2}([0,T];H)}^{2}\rightarrow 0.$
Thus we can pass to a further subsequence such that, $\mathbb{P}$-almost surely,
$\sup_{0\leqslant t\leqslant T}\left|\sqrt{\varepsilon}\int_{0}^{t}\langle\Lambda^{2\beta}\varphi(s),P_{m}\,\mathrm{d}W(s)\rangle\right|\rightarrow 0.$
Finally, Hölder’s inequality implies that
$\left|\int_{0}^{t}\left\langle P_{m}g-g,\Lambda^{2\beta}\varphi\right\rangle\,\mathrm{d}s\right|\lesssim\|\Lambda^{2\beta}\varphi\|_{L^{2}([0,T];H)}\left\|P_{m}g-g\right\|_{L^{2}([0,T];H)}\rightarrow 0.$
Therefore we conclude that, $\mathbb{P}$-almost surely, for all $t\in[0,T]$,
$\langle\bar{\theta}(t),\varphi(t)$ $\rangle+\int_{0}^{t}\left\langle\bar{\theta},\Lambda^{2\alpha}\varphi\right\rangle\,\mathrm{d}s=\left\langle\bar{\theta}(0),\varphi(0)\right\rangle+\int_{0}^{t}\langle\bar{\theta},\partial_{s}\varphi\rangle\,\mathrm{d}s$
$-\frac{1}{2}\int^{t}_{0}\left\langle R_{2}\bar{\theta},\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\bar{\theta}\right\rangle\,\mathrm{d}s+\frac{1}{2}\int_{0}^{t}\left\langle R_{1}\bar{\theta},\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\bar{\theta}\right\rangle\,\mathrm{d}s+\int_{0}^{t}\left\langle g,\Lambda^{2\beta}\varphi\right\rangle\,\mathrm{d}s.$
Step 2. $\bar{\theta}$ satisfies $H^{\alpha-2\beta}$-energy inequality.
Applying Itô’s formula to $\bar{\theta}_{\varepsilon,m}$, we deduce that $\mathbb{P}$-almost surely, for all $t\in[0,T]$,
$\frac{1}{2}$ $\left\|\bar{\theta}_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\bar{\theta}_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\left\|\bar{\theta}_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ (6.7)
$+\sqrt{\varepsilon}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle+\frac{\varepsilon}{2}\left\|P_{m}\Lambda^{\alpha}\right\|_{HS}^{2}t+\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m}(s),P_{m}g(s)\right\rangle\,\mathrm{d}s.$
Letting $\varepsilon\rightarrow 0$ in (6.7), it follows from the fact that $\bar{\theta}_{\varepsilon,m}\rightarrow\bar{\theta}\,\text{ in }\mathbb{X}_{\alpha,\beta}$ and the lower semi-continuity of the $H^{s}(\mathbb{T}^{2})$ norms that
$\frac{1}{2}\|\bar{\theta}(t)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\|\bar{\theta}(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\liminf_{\varepsilon\rightarrow 0}\frac{1}{2}\left\|\bar{\theta}_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\bar{\theta}_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s.$
Let $M^{\varepsilon,m}_{3}(t):=\int_{0}^{t}\sqrt{\varepsilon}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m}(s),P_{m}\,\mathrm{d}W(s)\right\rangle$. Applying Doob’s $L^{2}$ inequality, we obtain
$\mathbb{E}\left(\sup_{t\leqslant T}\left|M_{3}^{\varepsilon,m}(t)\right|^{2}\right)\leqslant 4\mathbb{E}\left\langle\langle M_{3}^{\varepsilon}\right\rangle\rangle_{T}\leqslant 4\varepsilon\mathbb{E}\int_{0}^{T}\left\|\bar{\theta}_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\rightarrow 0.$
Hence we can pass to a further subsequence so that $M_{3}^{\varepsilon,m}(t)\rightarrow 0$ almost surely. By the scaling regime (2.8), the correction term $\frac{\varepsilon}{2}\left\|P_{m}\Lambda^{\alpha}\right\|_{HS}^{2}t\rightarrow 0$. Finally, according to the weak convergence $\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m}\rightarrow\Lambda^{2\alpha-2\beta}\bar{\theta}$ in $L^{2}([0,T];H)$ and the strong convergence $P_{m}g\rightarrow g$ in $L^{2}([0,T];H)$, it holds that
$\left|\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta}_{\varepsilon,m},P_{m}g\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta},g\right\rangle\,\mathrm{d}s\right|\rightarrow 0.$
Therefore we conclude that, for all $t\in[0,T]$,
$\frac{1}{2}\left\|\bar{\theta}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\bar{\theta}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\left\|\bar{\theta}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\bar{\theta},g\right\rangle\,\mathrm{d}s.$
Step 3. $\bar{\theta}=\theta$ in $\mathbb{X}_{\alpha,\beta}.$
Thanks to the choice $\theta\in\mathcal{C}_{0}$, it follows from Definition 6.1 that $\theta=\bar{\theta}$ in $\mathbb{X}_{\alpha,\beta}$. Hence $\pi^{\varepsilon,m,\theta}$ converges to $\delta_{\theta}$ in law on $\mathbb{X}_{\alpha,\beta}$ as $\varepsilon\rightarrow 0$. ∎
We can now prove the large deviations lower bound restricted to $\mathcal{C}_{0}$.
Lemma 6.6 (Lower bound on $\mathcal{C}_{0}$).
For every $\varepsilon,\delta>0$ and $m\in\mathbb{N}_{+}$, let $\theta_{\varepsilon,\delta}$ be a stochastic generalized Leray solution of (1.3) in the sense of Definition 3.1 with initial data $\theta_{\varepsilon,\delta}(0)$ satisfying Assumption A, and let $\theta_{\varepsilon,m}$ be a solution of (3.7) with initial data $\theta_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$. Suppose that $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ is a weak-strong uniqueness class of (1.1).
(i) Let $\mu_{\varepsilon}=\mu_{\varepsilon,\delta}$ be the law of $\theta_{\varepsilon,\delta}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.7) holds for $(\varepsilon,\delta)$, then for any open set $G\subset\mathbb{X}_{\alpha,\beta}$,
$\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon}(G)\geqslant-\inf_{\theta\in G\cap\mathcal{C}_{0}}\mathcal{I}(\theta).$ (6.8)
(ii) Let $\mu_{\varepsilon}=\mu_{\varepsilon,m}$ be the law of $\theta_{\varepsilon,m}$ on $\mathbb{X}_{\alpha,\beta}$. Assume that the scaling regime (2.8) holds for $(\varepsilon,m)$, then for any open set $G\subset\mathbb{X}_{\alpha,\beta}$, the restricted lower bound (6.8) holds as well.
Proof.
(ii) is a direct consequence of Lemma 6.4, Lemma 6.5, and Lemma 2.5. And (i) can be proved in much the same way, the only difference being in the analysis of the pathwise $H^{\alpha-2\beta}$-energy inequality. More precisely, in the compactness analysis, we need to replace the Itô’s formula (6.7) by the $H^{\alpha-2\beta}$-energy inequality in Definition 3.1. ∎
Now we come to the proof of the restricted lower bound on $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$.
Proof of Proposition 6.3.
For any open set $G\subset\mathbb{X}_{\alpha,\beta}$ and any $\eta>0$, there exists an element $\theta^{0}\in G\cap\mathcal{C}$ such that
$\mathcal{I}(\theta^{0})\leqslant\inf_{\theta\in G\cap\mathcal{C}}\mathcal{I}(\theta)+\frac{\eta}{2}.$ (6.9)
By the definition of $\mathcal{C}$, there exists a sequence $\{\theta^{(n)}\}_{n\geqslant 1}\subset\mathcal{C}_{0}$, such that $\theta^{(n)}\rightarrow\theta^{0}$ in $\mathbb{X}_{\alpha,\beta}$ and $\mathcal{I}\left(\theta^{(n)}\right)\rightarrow\mathcal{I}(\theta^{0})$ as $n\rightarrow\infty$. It follows that
$\inf_{\theta\in G\cap\mathcal{C}_{0}}\mathcal{I}(\theta)\leqslant\mathcal{I}(\theta^{0})+\frac{\eta}{2}.$ (6.10)
Using (6.9), (6.10) and the arbitrariness of $\eta$, we obtain that
$\inf_{\theta\in G\cap\mathcal{C}_{0}}\mathcal{I}(\theta)=\inf_{\theta\in G\cap\mathcal{C}}\mathcal{I}(\theta).$
Hence (6.2) follows from (6.8).
7. Probabilistic Approaches to the Energy Equality
In this section, we are devoted to exploring the relationship between the deterministic energy equality and large deviations. For any $\alpha,\beta$ satisfying (1.2), we introduce the time-reversal operator $\mathfrak{T}_{T}:\mathbb{X}_{\alpha,\beta}\rightarrow\mathbb{X}_{\alpha,\beta}$, defined by $\mathfrak{T}_{T}\theta:=-\theta(T-\cdot)$.
Theorem 7.1.
Assume further that $\beta=\alpha/2.$ Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\alpha/2}$ be a weak-strong uniqueness class of (1.1) and $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$. Let $\mathcal{R}\subset\mathcal{C}$ satisfy $\mathcal{R}=\mathfrak{T}_{T}\mathcal{R}.$ Suppose that $\theta\in\mathcal{R}$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];H)$ in the sense of Definition 4.1. Then the kinetic energy equality holds:
$\frac{1}{2}\|\theta(T)\|_{H}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\|\theta(0)\|_{H}^{2}+\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s.$ (7.1)
Theorem 7.1 establishes a connection between the weak-strong uniqueness regularity class and the kinetic energy equality. However, the proof of this purely analytic result is based on large deviations for (3.7). The key argument relies on the time-reversibility property, and the proof will be provided later. Beyond time-reversibility, we emphasize that a Hamiltonian energy equality still holds in the case $\beta=\alpha/2+1/4$. This leads to the following strengthening of Theorem 7.1:
Theorem 7.1.
Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ be a weak-strong uniqueness class of (1.1) and $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$. Let $\mathcal{R}\subset\mathcal{C}$ satisfy $\mathcal{R}=\mathfrak{T}_{T}\mathcal{R}.$ Suppose that $\theta\in\mathcal{R}$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];H)$ in the sense of Definition 4.1. Then the $H^{\alpha-2\beta}$-energy equality holds:
$\frac{1}{2}\|\theta(T)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\|\theta(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\langle\Lambda^{2\alpha-2\beta}\theta,g\rangle\,\mathrm{d}s.$ (7.2)
The probabilistic method breaks down when $\beta\neq\alpha/2.$ Instead, we will provide an analytic proof for this more general case, following a similar approach to the one used in [GHW24, Theorem 2.7].
The remainder of this section will be dedicated to proving Theorem 7.1 and Theorem 7.1.
Lemma 7.2.
For any $\varepsilon>0$ and $m\in\mathbb{N}_{+}$, suppose that $\theta_{\varepsilon,m}$ is a solution of (3.7) with $\beta=\alpha/2$ and $\theta_{\varepsilon,m}(0)\sim\mathcal{G}(0,\varepsilon P_{m}/2)$. Then the law of $\theta_{\varepsilon,m}$ is invariant under $\mathfrak{T}$:
$\mathfrak{T}_{T}\theta_{\varepsilon,m}:=\left(-\theta_{\varepsilon,m}(T-t):0\leqslant t\leqslant T\right)\overset{\,\mathrm{d}}{=}\left(\theta_{\varepsilon,m}(t):0\leqslant t\leqslant T\right).$
Proof.
Let $\{e_{k}\}_{k\in\mathbb{Z}^{2}\backslash\{0\}}$ be the Fourier basis of $H$ introduced in Section 2, and we set $M=M(m):=\sum_{0<|k|\leqslant m}1.$ The solution $\theta_{\varepsilon,m}$ can be expanded as $\theta_{\varepsilon,m}=\sum_{0<|k|\leqslant m}X_{k}e_{k}$, leading to the expression $u_{\theta_{\varepsilon,m}}=R^{\perp}\theta_{\varepsilon,m}=\sum_{0<|k|\leqslant m}X_{k}R^{\perp}e_{k}.$ We denote by $u_{k}:=R^{\perp}e_{k}$ and $B_{k,l,n}:=\langle u_{n}\cdot\nabla e_{l},e_{k}\rangle$, then (3.7) is equivalent to the following stochastic differential equations on $\mathbb{R}^{M}:$
[…]
Let $\mathcal{G}$ denote the Gaussian measure on $\mathbb{R}^{M}$ with mean zero and covariance $\varepsilon I/2$, where $I$ is the identity operator. Let $\mathcal{L}$ be the infinitesimal generator of the Markov process $X=(X_{k})_{0<|k|\leqslant m}$, decomposed as $\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{2}$, where
$\mathcal{L}_{1}f(x)$ $=\sum_{0<|k|\leqslant m}\left(-|k|^{2\alpha}x_{k}\partial_{k}f(x)+\frac{\varepsilon}{2}|k|^{2\alpha}\partial_{k}^{2}f(x)\right),$
$\mathcal{L}_{2}f(x)$ $=-\sum_{0<|k|\leqslant m}\sum_{0<|l|\leqslant m}\sum_{0<|n|\leqslant m}B_{k,l,n}x_{l}x_{n}\partial_{k}f(x),\quad\text{for all }f\in C_{b}^{2}\left(\mathbb{R}^{M}\right).$
For any $f,g\in C_{b}^{2}\left(\mathbb{R}^{M}\right)$, it follows from the integration by parts formula that
$\left\langle f,\mathcal{L}_{1}g\right\rangle_{L^{2}(\mathcal{G})}=\left\langle\mathcal{L}_{1}f,g\right\rangle_{L^{2}(\mathcal{G})},\quad\langle f,\mathcal{L}_{2}g\rangle_{L^{2}(\mathcal{G})}=-\langle\mathcal{L}_{2}f,g\rangle_{L^{2}(\mathcal{G})},$
which implies that $\mathcal{G}$ is an invariant measure of (7.3). The infinitesimal generator of $-X(\cdot)$ is $\mathcal{L}_{1}-\mathcal{L}_{2}$ and the infinitesimal generator of $X(T-\cdot)$ is $\mathcal{L}^{*}=\mathcal{L}_{1}^{*}+\mathcal{L}_{2}^{*}=\mathcal{L}_{1}-\mathcal{L}_{2}$, where denotes the adjoint operator in $L^{2}(\mathcal{G})$. This demonstrates the time-reversibility $\mathfrak{T}_{T}\theta_{\varepsilon,m}\overset{\mathrm{d}}{=}\theta_{\varepsilon,m}$.
Proof of Theorem 7.1.
We begin by proving that $\mathcal{I}(\mathfrak{T}_{T}\theta)=\mathcal{I}(\theta).$ Suppose that $\varepsilon,m$ and $\theta_{\varepsilon,m}$ satisfy the conditions in Lemma 7.2. Let $\mu_{\varepsilon,m}$ be the law of $\theta_{\varepsilon,m}$ on $\mathbb{X}_{\alpha,\alpha/2}$ and let $\tilde{\mu}_{\varepsilon,m}:=\mu_{\varepsilon,m}\circ\mathfrak{T}_{T}^{-1}$. For any $\eta>0,$ it follows from Lemma 4.4 that there exists an open set $G$ in $\mathbb{X}_{\alpha,\alpha/2}$ such that $\theta\in G$ and $\mathcal{I}(\hat{\theta})>\mathcal{I}(\mathfrak{T}_{T}\theta)-\eta$ for all $\hat{\theta}\in\mathfrak{T}_{T}\overline{G}.$ Moreover, the continuity of $\mathfrak{T}_{T}:\mathbb{X}_{\alpha,\alpha/2}\rightarrow\mathbb{X}_{\alpha,\alpha/2}$ implies that $\mathfrak{T}_{T}^{-1}\overline{G}=\mathfrak{T}_{T}\overline{G}$ is closed in $\mathbb{X}_{\alpha,\alpha/2}.$ According to Proposition 6.3, Lemma 7.2 and Proposition 5.1, we have
$-\mathcal{I}(\theta)\leqslant\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)$ $=\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\tilde{\mu}_{\varepsilon,m}(G)$
$\leqslant\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\tilde{\mu}_{\varepsilon,m}(\overline{G})$
$=\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(\mathfrak{T}_{T}\overline{G})$
$\leqslant-\inf_{\hat{\theta}\in\mathfrak{T}_{T}\overline{G}}\mathcal{I}(\hat{\theta})\leqslant-\mathcal{I}(\mathfrak{T}_{T}\theta)+\eta.$
Due to the arbitrariness of $\eta,$ we deduce that $\mathcal{I}(\mathfrak{T}_{T}\theta)\leqslant\mathcal{I}(\theta).$ Since $\mathcal{R}=\mathfrak{T}_{T}\mathcal{R}$, we can replace $\theta$ by $\mathfrak{T}_{T}\theta$ to obtain that $\mathcal{I}(\mathfrak{T}_{T}\theta)=\mathcal{I}(\theta).$
Thanks to Remark 4.3, $g$ is the unique element in $L^{2}([0,T];H)$ such that $\theta$ is a weak solution of (1.1) for $g$. Thus
$\mathcal{I}(\theta)=\|\theta(0)\|^{2}_{H}+\frac{1}{2}\big{\|}g\big{\|}^{2}_{L^{2}([0,T];H)}.$ (7.4)
Now we define $\tilde{g}:[0,T]\times\mathbb{T}^{2}\rightarrow\mathbb{R}$ by $\tilde{g}(t,x):=g(T-t,x)-2\Lambda^{\alpha}(\mathfrak{T}_{T}\theta)$. We have $\tilde{g}\in L^{2}([0,T];H)$ since $\theta\in\mathcal{R}\subset\mathbb{X}_{\alpha,\alpha/2}.$ Moreover, a direct calculation shows that $\tilde{\theta}$ is a weak solution of (1.1) with $\tilde{g}$ in the sense of Definition 4.1. Hence
$\mathcal{I}(\mathfrak{T}_{T}\theta)$ $=\|\theta(T)\|^{2}_{H}+\frac{1}{2}\big{\|}\tilde{g}\big{\|}^{2}_{L^{2}([0,T];H)}$ (7.5)
$=\|\theta(T)\|^{2}_{H}+\frac{1}{2}\big{\|}g\big{\|}^{2}_{L^{2}([0,T];H)}+2\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}-2\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s.$
Consequently, in combination with $\mathcal{I}(\mathfrak{T}_{T}\theta)=\mathcal{I}(\theta)$, (7.4) and (7.5) lead to the kinetic energy equality
$\frac{1}{2}\|\theta(T)\|_{H}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\|\theta(0)\|_{H}^{2}+\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s.$
Corollary 7.3.
Let $\varepsilon>0$ and $m\in\mathbb{N}_{+}$ satisfy the scaling regime (2.8) and let $\theta_{\varepsilon,m},\mu_{\varepsilon,m}$ be as in Proposition 6.3. Suppose that $\theta\in\mathbb{X}_{\alpha,\alpha/2}$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];H)$ in the sense of Definition 4.1.
(i) Assume that
$\frac{1}{2}\|\theta(T)\|_{H}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}>\frac{1}{2}\|\theta(0)\|_{H}^{2}+\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s,$
then $\theta$ violates the large deviations lower bound:
$\inf_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2}:\,\theta\in G\}}\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)<-\mathcal{I}(\theta).$
(ii) Assume that
$\frac{1}{2}\|\theta(T)\|_{H}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}<\frac{1}{2}\|\theta(0)\|_{H}^{2}+\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s,$
then $\mathfrak{T}_{T}\theta$ violates the large deviations lower bound:
$\inf_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2}:\,\mathfrak{T}_{T}\theta\in G\}}\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)<-\mathcal{I}(\mathfrak{T}_{T}\theta).$
Proof.
We first prove (i). Assume the converse, specifically that
$\inf_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2}:\,\theta\in G\}}\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)\geqslant-\mathcal{I}(\theta).$ (7.6)
We will show that $\theta$ satisfies the kinetic energy inequality
$\frac{1}{2}\|\theta(T)\|_{H}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\leqslant\frac{1}{2}\|\theta(0)\|_{H}^{2}+\int_{0}^{T}\langle\Lambda^{\alpha}\theta,g\rangle\,\mathrm{d}s,$ (7.7)
and thereby obtain a contradiction. Let $\tilde{\mu}_{\varepsilon,m}:=\mu_{\varepsilon,m}\circ\mathfrak{T}_{T}^{-1}$. Using (7.6), Lemma 7.2 and Proposition 5.1, for any open set $G\subset\mathbb{X}_{\alpha,\alpha/2}$,
$-\mathcal{I}(\theta)\leqslant\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(G)$ $=\liminf_{\varepsilon\rightarrow 0}\varepsilon\log\tilde{\mu}_{\varepsilon,m}(G)$
$\leqslant\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\tilde{\mu}_{\varepsilon,m}(\overline{G})$
$=\limsup_{\varepsilon\rightarrow 0}\varepsilon\log\mu_{\varepsilon,m}(\mathfrak{T}_{T}\overline{G})$
$\leqslant-\inf_{\hat{\theta}\in\overline{G}}\mathcal{I}(\mathfrak{T}_{T}\hat{\theta}).$
Since $\mathcal{I}$ is lower semi-continuous, $\sup_{\{G\text{ is open in }\mathbb{X}_{\alpha,\alpha/2}:\,\theta\in G\}}\inf_{\hat{\theta}\in\overline{G}}\mathcal{I}(\mathfrak{T}_{T}\hat{\theta})=\mathcal{I}(\mathfrak{T}_{T}\theta).$ Therefore, $\mathcal{I}(\theta)\geqslant\mathcal{I}(\mathfrak{T}_{T}\theta)$. Then it follows from (7.4) and (7.5) that (7.7) holds. We can apply this argument again, with $\theta$ replaced by $\mathfrak{T}_{T}\theta$, to obtain (ii). ∎
The following lemma will be utilized in the proof of Theorem 7.1.
Lemma 7.4.
Let $\mathcal{C}_{0}\subset\mathbb{X}_{\alpha,\beta}$ be a weak-strong uniqueness class of (1.1) and let $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$. Suppose that $\theta\in\mathcal{C}$ is a weak solution of (1.1) for some $g\in L^{2}([0,T];H)$ in the sense of Definition 4.1. Then the $H^{\alpha-2\beta}$-energy inequality holds:
$\frac{1}{2}\|\theta(T)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\|\theta(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\langle\Lambda^{2\alpha-2\beta}\theta,g\rangle\,\mathrm{d}s.$
Proof.
By the definition of $\mathcal{C}$, there exists a sequence $\{\theta^{(n)}\}\subset\mathcal{C}_{0}$, such that $\theta^{(n)}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\beta}$ and $\mathcal{I}\big{(}\theta^{(n)}\big{)}\rightarrow\mathcal{I}(\theta)$. Let $g^{(n)}\in L^{2}([0,T];H)$ be the control such that $\theta^{(n)}$ solves (1.1) weakly. We claim that $g^{(n)}\rightarrow g$ in $L^{2}([0,T];H)$. In fact, $\mathcal{I}\big{(}\theta^{(n)}\big{)}\rightarrow\mathcal{I}(\theta)$ implies that $\{g^{(n)}\}$ is uniformly bounded in $L^{2}([0,T];H),$ hence weakly compact. For any $\varphi\in C^{\infty}([0,T];H^{\infty}(\mathbb{T}^{2}))$, let $\Lambda_{1}^{T}(\varphi,\cdot),F^{T}(\varphi,\cdot)$ be defined as in Section 4. Suppose that $g_{0}$ is a limit point of $\{g^{(n)}\}$ with respect to the weak topology. Then it follows from (4.14) that we can let $n\rightarrow\infty$ in (4.13) with $\Psi^{\theta}$ replaced by $\Psi^{\theta^{(n)}}$ to obtain that
$F^{T}(\varphi,\theta)=\langle g_{0},\Lambda^{2\beta}\varphi\rangle_{L^{2}([0,T];H)}.$
That is, $\theta$ is a weak solution of (1.1) for the control $g_{0}.$ We deduce from Remark 4.3 that $g=g_{0}$ in $L^{2}([0,T];H)$. Therefore, $g^{(n)}\rightarrow g$ weakly in $L^{2}([0,T];H)$. Since $\mathcal{I}\big{(}\theta^{(n)}\big{)}\rightarrow\mathcal{I}(\theta)$ implies $\|g^{(n)}\|_{L^{2}([0,T];H)}\rightarrow\|g\|_{L^{2}([0,T];H)},$ we can conclude that $g^{(n)}\rightarrow g$ (strongly) in $L^{2}([0,T];H).$
As in the proof of Lemma 6.5, $\theta^{(n)}\in\mathcal{C}_{0}$ satisfies the $H^{\alpha-2\beta}$-energy inequality:
$\frac{1}{2}\|\theta^{(n)}(T)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\|\theta^{(n)}(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\|\theta^{(n)}(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\langle\Lambda^{2\alpha-2\beta}\theta^{(n)},g^{(n)}\rangle\,\mathrm{d}s.$
Letting $n\rightarrow\infty,$ due to the strong convergence of $g^{(n)}$, a repetition of Step 2 in the proof of Lemma 6.5 yields that
$\frac{1}{2}\|\theta(T)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\|\theta(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\langle\Lambda^{2\alpha-2\beta}\theta,g\rangle\,\mathrm{d}s.$
Proof of Theorem 7.1. We are now in the setting of Lemma 7.4. It follows that
$\frac{1}{2}\|\theta(T)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\|\theta(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\|\theta(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\langle\Lambda^{2\alpha-2\beta}\theta,g\rangle\,\mathrm{d}s.$ (7.8)
Recall that the map $\tilde{g}$ defined by $\tilde{g}(t,x):=g(T-t,x)-2\Lambda^{\alpha}(\mathfrak{T}_{T}\theta)$ is an element of $L^{2}([0,T];H)$ and $\mathfrak{T}_{T}\theta$ is a weak solution of (1.1) for $\tilde{g}$. Since $\mathfrak{T}_{T}\theta\in\mathcal{R}\subset\mathcal{C}$, we can apply Lemma 7.4 for $\mathfrak{T}_{T}\theta$ and $\tilde{g}$ to obtain
$\frac{1}{2}\|\theta(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ $+\int_{0}^{T}\|\theta(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\|\theta(T)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\langle\Lambda^{2\alpha-2\beta}\mathfrak{T}_{T}\theta,\tilde{g}\rangle\,\mathrm{d}s$ (7.9)
$=\frac{1}{2}\|\theta(T)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}-\int_{0}^{T}\langle\Lambda^{2\alpha-2\beta}\theta,g\rangle\,\mathrm{d}s+2\int_{0}^{T}\|\theta(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s.$
Combining (7.8) with (7.9), we conclude that (7.2) holds.
8. Characterization of Quasi-potential
In this section, we will analyze the explicit representation of the quasi-potential $\mathcal{U}.$ This potentially indicates a new probabilistic perspective of the uniqueness problem for the deterministic SQG equation.
8.1. Conditional equivalence
To begin with, we need to extend the definition of the weak solution to infinite time intervals.
Definition 8.1.
Let $g\in L_{loc}^{2}([0,\infty);H)$. We say that $\theta$ is a weak solution of
$\partial_{t}\theta=-\Lambda^{2\alpha}\theta-R^{\perp}\theta\cdot\nabla\theta+\Lambda^{2\beta}g,$ (8.1)
if $\theta\in L^{\infty}_{loc}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2}))\cap L^{2}_{loc}([0,\infty);H^{2\alpha-2\beta}(\mathbb{T}^{2}))\cap C([0,\infty);H_{w}^{\alpha-2\beta}(\mathbb{T}^{2})),$ and for every $\varphi\in C^{\infty}([0,\infty);H^{\infty}(\mathbb{T}^{2})),$ for all $t>0,$
$\langle\theta(t),\varphi(t)\rangle$ $=\langle\theta(0),\varphi(0)\rangle+\int^{t}_{0}\langle\theta,\partial_{s}\varphi\rangle\,\mathrm{d}s-\int^{t}_{0}\langle\theta,\Lambda^{2\alpha}\varphi\rangle\,\mathrm{d}s$ (8.2)
$-\frac{1}{2}\int^{t}_{0}\left\langle R_{2}\theta,\left[\Lambda,\partial_{1}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}s+\frac{1}{2}\int_{0}^{t}\left\langle R_{1}\theta,\left[\Lambda,\partial_{2}\varphi\right]\Lambda^{-1}\theta\right\rangle\,\mathrm{d}s+\int^{t}_{0}\langle\Lambda^{2\beta}\varphi,g\rangle\,\mathrm{d}s.$
With the above definition, we can similarly define the weak-strong uniqueness class on $[0,\infty)$ as in Definition 6.1. Let
$\mathbb{X}^{\infty}_{\alpha,\beta}:=L^{2}_{loc}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2}))\cap L^{2}_{w,loc}([0,\infty);H^{2\alpha-2\beta}(\mathbb{T}^{2}))\cap C([0,\infty);H^{\alpha-2\beta}_{w}(\mathbb{T}^{2})).$
Definition 8.2 (Weak-Strong Uniqueness Class on $[0,\infty)$).
We say $\mathcal{C}_{0}\subset\mathbb{X}^{\infty}_{\alpha,\beta}$ is a weak-strong uniqueness class of (8.1) for every $g\in L^{2}_{loc}([0,\infty);H)$ if, for arbitrary two weak solutions $\theta_{1},\theta_{2}$ of the skeleton equation (8.1) in the sense of Definition 8.1 with the same initial data $\theta_{1}(0)=\theta_{2}(0)$ and the same control $g$, we have $\theta_{1}=\theta_{2}$ in $\mathbb{X}_{\alpha,\beta}^{\infty}$ as long as $\theta_{1}\in\mathcal{C}_{0}$ and $\theta_{2}$ satisfies the $H^{\alpha-2\beta}$-energy inequality: for every $t>0,$
$\frac{1}{2}\|\theta_{2}(t)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\|\theta_{2}(s)\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\left\|\theta_{2}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\langle\Lambda^{2\alpha-2\beta}\theta_{2},g\rangle\,\mathrm{d}s.$ (8.3)
Let $\mathcal{C}_{0}$ be a weak-strong uniqueness class on $[0,\infty).$ Denote
$\mathcal{A}_{0}:=\big{\{}\theta\in\mathcal{C}_{0}:\partial_{t}\theta\in L^{2}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2})),R^{\perp}\theta\cdot\nabla\theta\in L^{2}([0,\infty);H^{\alpha-2\beta}(\mathbb{T}^{2}))$ ,
[…]
For any $\tilde{\theta}:(-\infty,0]\times\mathbb{T}^{2}\rightarrow\mathbb{R}$ that satisfying $-\tilde{\theta}(-\cdot)\in\mathcal{A}_{0}$, we define
$S_{-\infty}(\tilde{\theta}):=\frac{1}{2}\int_{-\infty}^{0}\|\partial_{t}\tilde{\theta}(s)+\Lambda^{2\alpha}\tilde{\theta}(s)+R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s)\|_{H^{-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s,$
and define the map $\mathcal{U}:H^{\alpha-2\beta}(\mathbb{T}^{2})\rightarrow[0,\infty)$ as
$\mathcal{U}(\phi):=\inf\left\{S_{-\infty}(\tilde{\theta}):-\tilde{\theta}(-\cdot)\in\mathcal{A}_{0},\,\tilde{\theta}(0)=\phi\right\}.$ (8.4)
Proposition 8.3.
Let $\mathcal{C}_{0},\mathcal{A}_{0},\mathcal{U}$ and $S_{-\infty}$ be defined as above.
(i) For every $\phi\in H^{\alpha-2\beta}(\mathbb{T}^{2})$, it holds that $\mathcal{U}(\phi)\geqslant\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$.
(ii) The equality $\mathcal{U}(\phi)=\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ holds for a given $\phi\in H^{\alpha-2\beta}(\mathbb{T}^{2})$ as long as the equation
$\left\{\begin{array}[]{l}\partial_{t}\bar{\theta}=-\Lambda^{2\alpha}\bar{\theta}-R^{\perp}\bar{\theta}\cdot\nabla\bar{\theta},\\ \bar{\theta}(0)=\phi,\end{array}\right.$ (8.5)
admits a unique weak solution in $\mathcal{A}_{0}$ in the sense of Definition 8.1.
Proof.
(i). It is sufficient to discuss the case that $\mathcal{U}(\phi)<\infty.$ In this case, there exists a function $\tilde{\theta}:(-\infty,0]\times\mathbb{T}^{2}\rightarrow\mathbb{R}$ such that $-\tilde{\theta}(-\cdot)\in\mathcal{A}_{0}$, $\tilde{\theta}(0)=\phi$ and $\langle R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s),\tilde{\theta}(s)\rangle_{H^{\alpha-2\beta}(\mathbb{T}^{2})}=0$ for all $s<0.$ Then it follows from the definition of $\mathcal{A}_{0}$ that
$\left\|\Lambda^{-2\beta}(\partial_{t}\tilde{\theta}(s)+\Lambda^{2\alpha}\tilde{\theta}(s)+R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s))\right\|_{H}^{2}$
$=$ $\left\|\Lambda^{-2\beta}(\partial_{t}\tilde{\theta}(s)-\Lambda^{2\alpha}\tilde{\theta}(s)+R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s))\right\|_{H}^{2}+4\|\Lambda^{2\alpha-2\beta}\tilde{\theta}(s)\|_{H}^{2}$
$+4\left\langle\Lambda^{-2\beta}(\partial_{t}\tilde{\theta}(s)-\Lambda^{2\alpha}\tilde{\theta}(s)+R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s)),\Lambda^{2\alpha-2\beta}\tilde{\theta}(s)\right\rangle$
$=$ $\left\|\Lambda^{-2\beta}(\partial_{t}\tilde{\theta}(s)-\Lambda^{2\alpha}\tilde{\theta}(s)+R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s))\right\|_{H}^{2}+4\left\langle\Lambda^{\alpha-2\beta}\partial_{t}\tilde{\theta}(s),\Lambda^{\alpha-2\beta}\tilde{\theta}(s)\right\rangle.$
Therefore, by the definition of $S_{-\infty}$, we have
$S_{-\infty}(\tilde{\theta})$ $=\frac{1}{2}\int_{-\infty}^{0}\left\|\Lambda^{-2\beta}(\partial_{t}\tilde{\theta}(s)-\Lambda^{2\alpha}\tilde{\theta}(s)+R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s))\right\|_{H}^{2}\,\mathrm{d}s+\int_{-\infty}^{0}\frac{\mathrm{d}}{\mathrm{d}s}\|\tilde{\theta}(s)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$
$=\frac{1}{2}\int_{-\infty}^{0}\left\|\Lambda^{-2\beta}(\partial_{t}\tilde{\theta}(s)-\Lambda^{2\alpha}\tilde{\theta}(s)+R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s))\right\|_{H}^{2}\,\mathrm{d}s+\|\tilde{\theta}(0)\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}.$ (8.6)
Thus $\tilde{\theta}(0)=\phi$ implies $\mathcal{U}(\phi)\geqslant\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$.
(ii). Now we aim to prove that the equality $\mathcal{U}(\phi)=\|\phi\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ holds as long as (8.5) admits a weak solution in $\mathcal{A}_{0}.$ Assume that $\bar{\theta}$ is the solution. Then
$\tilde{\theta}(\cdot):=-\bar{\theta}(-\cdot)$
satisfies $\partial_{t}\tilde{\theta}(s)-\Lambda^{2\alpha}\tilde{\theta}(s)+R^{\perp}\tilde{\theta}(s)\cdot\nabla\tilde{\theta}(s)=0$ for all $s\leqslant 0$ and hence the first term on the right-hand side of (8.1) is 00. This completes the proof. ∎
Remark 8.4.
We provide further remarks on the choice of $\mathcal{A}_{0}$. It is evident that such a conditional equivalence does not rely on the concrete concept of ’weak-strong uniqueness’ in the definition of $\mathcal{A}_{0}$. In other words, the conditional equivalence between the restricted quasi-potential and the Gaussian rate function depends only on the existence of solutions of (8.5) within the restriction class of the quasi-potential. Therefore, the class $\mathcal{A}_{0}$ can be replaced by other regularity classes.
Nevertheless, to demonstrate the relationships between the non-Gaussian large deviations (see Definition 8.5 below) in the equilibrium of the stochastic PDE (1.3) and the uniqueness problem of the PDE (8.5), one needs to specify the restriction class of the quasi-potential to govern the equilibrium-large deviations rate. As indicated in [CP22, Proposition 4.1], this restriction class depends on where the upper and lower bounds of the uniform dynamical large deviations match. To invoke the dynamical large deviations result in this paper, one expects to define $\mathcal{A}_{0}$ with trajectories lying in the $\mathcal{I}$-closure class $\mathcal{C}$, which is the domain where the upper and lower bounds match.
However, the extension of the lower bound from the restriction class $\mathcal{C}_{0}$ to $\mathcal{C}$ technically relies on the choice of the random initial data. If the random initial data was replaced by a deterministic initial data, it is unclear whether the extension of the lower bound holds on $\mathcal{C}$, as well as the uniform dynamical large deviations lower bound. Therefore, we specify that the trajectories of $\mathcal{A}_{0}$ lie only in the smaller set $\mathcal{C}_{0}$ instead of $\mathcal{C}$.
8.2. Potential relationships to open problems for PDEs
To demonstrate relationships to open problems for PDEs, we introduce the concept of non-Gaussian large deviations in equilibrium for stochastic PDE (1.3).
Definition 8.5 (Non-Gaussian large deviations).
Let $\{\nu_{\varepsilon}\}_{\varepsilon>0}$ be a family of probability measures on $H_{w}^{\alpha-2\beta}(\mathbb{T}^{2})$. We say that $\{\nu_{\varepsilon}\}_{\varepsilon>0}$ satisfies non-Gaussian large deviations on $H_{w}^{\alpha-2\beta}(\mathbb{T}^{2})$, if it satisfies large deviations on $H_{w}^{\alpha-2\beta}(\mathbb{T}^{2})$ with speed $\varepsilon^{-1}$ and rate function $I_{0}(\cdot)\neq\|\cdot\|^{2}_{H^{\alpha-2\beta}(\mathbb{T}^{2})}.$
Let $\{\nu_{\varepsilon,\delta}\}_{\varepsilon,\delta>0}$ be a family of ergodic invariant measures of (1.3). Inspired by [BC17] and [CP22], the large deviations for $\{\nu_{\varepsilon,\delta}\}_{\varepsilon,\delta>0}$ under a scaling regime $(\varepsilon,\delta(\varepsilon))\rightarrow(0,0)$ strongly rely on uniform dynamical large deviations, and its rate function is expected to be governed by the quasi-potential. Based on the discussion in Remark 8.4, one can only expect to observe the rate function for the lower bound governed by the restricted quasi-potential (8.4). This connects to the open problem of uniqueness for the SQG equation (8.5).
As long as one can prove large deviations for $\{\nu_{\varepsilon,\delta(\varepsilon)}\}_{\varepsilon>0}$ with rate function (8.4), then Proposition 8.3 indicates the following relationships. On the one hand, if a solution of (8.5) exists in $\mathcal{A}_{0}$ (which implies uniqueness), then, by applying Proposition 8.3, one can observe that the expected rate function (8.4) equals to the Gaussian rate function. Conversely, if non-Gaussian large deviations can be proven for $\{\nu_{\varepsilon,\delta(\varepsilon)}\}_{\varepsilon>0}$ in the sense of Definition 8.5, the uniqueness problem of (8.5) could be answered negatively. Therefore, Proposition 8.3 suggests a new perspective for understanding the uniqueness problem of the PDE (8.5) through probabilistic approaches. However, rigorously establishing this relationship requires proving large deviations in equilibrium, which lies beyond the scope of this paper.
Remark 8.6.
It is worth pointing out that in [Tot20], the existence of Gaussian stationary solutions for (1.3) with $\delta=0$ is proved using the energy solutions approach. However, this does not contradict the discussion of non-Gaussian large deviations for invariant measures in the sense of Definition 8.5, as the aforementioned non-Gaussian large deviations focus on fluctuations in spaces with higher regularity due to the regularization of the noise and the choice of the scaling regime $(\varepsilon,\delta(\varepsilon))$.
On the other hand, in [HLZZ24], it was shown that there exist infinitely many non-Gaussian ergodic stationary solutions of (1.3) with $\delta=0$, which lie in the Besov space $B^{-1/2}_{p,1}$ for any $p\geqslant 2.$ In their work, the parameters $\alpha$ and $\beta$ satisfy $\alpha\in[0,3/4)$ and $\beta\in[0,1/8).$ We hope to bridge the discussion between non-Gaussian ergodic stationary solutions and the uniqueness problem of the PDE (8.5) via the non-Gaussian large deviations.
Appendix A A Examples of the Weak-strong Uniqueness Classes
In this section, we provide some examples of the weak-strong uniqueness classes $\mathcal{C}_{0}$ and their $\mathcal{I}$-closure $\mathcal{C}=\overline{\mathcal{C}_{0}}^{\mathcal{I}}$ for different choices of $\alpha$ and $\beta$ satisfying (1.2). These examples cover subcritical, critical and supercritical cases.
Example A.1.
Suppose that $11/2<\alpha<$ 1 and $\beta=\alpha/2+1/4$. Let
$\mathcal{C}_{0}^{1}=\{\theta\in\mathbb{X}_{\alpha,\alpha/2+1/4}:\theta\in L^{r}\left([0,T];L^{p}(\mathbb{T}^{2})\right),\,\frac{1}{p}+\frac{\alpha}{r}=\alpha-\frac{1}{2},\,p\geqslant 1,\,r>0\}.$
For every $g\in L^{2}([0,T];H)$, assume that $\theta_{1}\in\mathcal{C}_{0}^{1}$ and $\theta_{2}\in\mathbb{X}_{\alpha,\alpha/2+1/4}$ solve the skeleton equation (1.1) in the sense of Definition 4.1 with the same initial data $\theta_{1}(0)=\theta_{2}(0)\in H^{-1/2}(\mathbb{T}^{2})$, and suppose further that $\theta_{2}$ satisfy the $H^{-1/2}$-energy inequality: for every $t\in[0,T]$,
$\frac{1}{2}\|\theta_{2}(t)\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\|\theta_{2}(s)\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\left\|\theta_{2}(0)\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\langle\Lambda^{\alpha-1}\theta_{2},g\rangle\,\mathrm{d}s.$ (A.1)
Then $\theta_{1}=\theta_{2}$ in $\mathbb{X}_{\alpha,\alpha/2+1/4}$.
Proof.
The same conclusion holds in the case where $\beta=\alpha/2$, which was stated in [CW99, Theorem 2.2, Remark 2.3]. For every $t\in[0,T]$, we expand $\|\theta_{1}(t)-\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}$ as
$\|\theta_{1}(t)-\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}=2\langle\Lambda^{-1/2}\theta_{1}(t),\Lambda^{-1/2}\theta_{2}(t)\rangle+\|\theta_{1}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}+\|\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}$ (A.2)
and estimate these three terms respectively. For the first term of the right-hand side of (A.2), we need the following equality:
$\quad\int_{0}^{t}\left(2\langle\Lambda^{\alpha-1/2}\theta_{1},\Lambda^{\alpha-1/2}\theta_{2}\rangle+\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}\theta_{1}\rangle+\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}\theta_{2}\rangle\right)\,\mathrm{d}s$ (A.3)
$=-\langle\Lambda^{-1/2}\theta_{2}(t),\Lambda^{-1/2}\theta_{1}(t)\rangle+\langle\Lambda^{-1/2}\theta_{2}(0),\Lambda^{-1/2}\theta_{1}(0)\rangle+\int_{0}^{t}\langle\Lambda^{\alpha-1/2}(\theta_{2}+\theta_{1}),g\rangle\,\mathrm{d}s.$
For any $\epsilon>0$, let $\eta_{\epsilon}$ be the standard convolution kernel on $\mathbb{T}^{2}.$ For any real-valued function $f$, we define $f^{(n)}$ by $f^{(n)}:=f\ast\eta_{1/n}$. Since $\Lambda^{-1}\theta^{(n)}_{i}$ can be taken as a test function for all $n\geqslant 1$ and $i=1,2,$ it follows that
$\quad\int_{0}^{t}\left(-\langle\Lambda^{-1/2}\theta_{2},\partial_{t}\Lambda^{-1/2}\theta_{1}^{(n)}\rangle+\langle\Lambda^{\alpha-1/2}\theta_{2},\Lambda^{\alpha-1/2}\theta_{1}^{(n)}\rangle+\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}\theta_{1}^{(n)}\rangle\right)\,\mathrm{d}s$
$=-\langle\Lambda^{-1/2}\theta_{2}(t),\Lambda^{-1/2}\theta_{1}^{(n)}(t)\rangle+\langle\Lambda^{-1/2}\theta_{2}(0),\Lambda^{-1/2}\theta_{1}^{(n)}(0)\rangle+\int_{0}^{t}\langle\Lambda^{\alpha-1/2}\theta_{2},g\rangle\,\mathrm{d}s,$
and
$\quad\int_{0}^{t}\left(-\langle\Lambda^{-1/2}\theta_{1},\partial_{t}\Lambda^{-1/2}\theta_{2}^{(n)}\rangle+\langle\Lambda^{\alpha-1/2}\theta_{1},\Lambda^{\alpha-1/2}\theta_{2}^{(n)}\rangle+\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}\theta_{2}^{(n)}\rangle\right)\,\mathrm{d}s$
$=-\langle\Lambda^{-1/2}\theta_{1}(t),\Lambda^{-1/2}\theta_{2}^{(n)}(t)\rangle+\langle\Lambda^{-1/2}\theta_{1}(0),\Lambda^{-1/2}\theta_{2}^{(n)}(0)\rangle+\int_{0}^{t}\langle\Lambda^{\alpha-1/2}\theta_{1},g\rangle\,\mathrm{d}s.$
It suffices to pass to the limit $n\rightarrow\infty$ to get (A.3). Using Fubini’s theorem and the symmetry of the convolution kernel, we have
$\int_{0}^{t}$ $\langle\Lambda^{-1/2}\theta_{2},\partial_{t}\Lambda^{-1/2}\theta_{1}^{(n)}\rangle\,\mathrm{d}s+\int_{0}^{t}\langle\Lambda^{-1/2}\theta_{1},\partial_{t}\Lambda^{-1/2}\theta_{2}^{(n)}\rangle\,\mathrm{d}s$
$=\langle\Lambda^{-1/2}\theta_{2}(t),\Lambda^{-1/2}\theta_{1}^{(n)}(t)\rangle-\langle\Lambda^{-1/2}\theta_{2}(0),\Lambda^{-1/2}\theta_{1}^{(n)}(0)\rangle.$
Now we focus on the convergence of the nonlinear terms. Since $\theta_{1}\in\mathcal{C}_{0}^{1}$,
$\quad\left|\int_{0}^{t}\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}(\theta_{1}-\theta^{(n)}_{1})\rangle\,\mathrm{d}s\right|$ (A.4)
$\lesssim\int_{0}^{t}\|\theta_{2}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}\|\nabla\Lambda^{-1}(\theta_{1}-\theta^{(n)}_{1})\|_{L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})}\|u_{\theta_{2}}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2};\mathbb{R}^{2})}\,\mathrm{d}s$
$\lesssim\int_{0}^{t}\|\theta_{1}-\theta^{(n)}_{1}\|_{L^{p}(\mathbb{T}^{2})}\|\theta_{2}\|_{H^{1/p}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$
$\lesssim\int_{0}^{t}\|\theta_{1}-\theta^{(n)}_{1}\|_{L^{p}(\mathbb{T}^{2})}\|\theta_{2}\|_{H^{-1/2}(\mathbb{T}^{2})}^{2/r}\left\|\theta_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2(1-1/r)}\,\mathrm{d}s$
$\lesssim\left(\int_{0}^{t}\|\theta_{1}-\theta^{(n)}_{1}\|_{L^{p}(\mathbb{T}^{2})}^{r}\left\|\theta_{2}\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)^{1/r}\left(\int_{0}^{t}\left\|\theta_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)^{1-1/r}$
$\lesssim\sup_{t\in[0,T]}\left\|\theta_{2}(t)\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{2/r}\|\theta_{1}-\theta^{(n)}_{1}\|_{L^{r}\left([0,T];L^{p}(\mathbb{T}^{2})\right)}\left(\int_{0}^{t}\left\|\theta_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)^{1-1/r}\rightarrow 0.$
Here we apply Hölder’s inequality in the second line, Calderón-Zygmund inequality and Sobolev embedding theorem in the third line, Gagliardo-Nirenberg interpolation inequality in the fourth line and Hölder’s inequality in the fifth line. Similarly,
$\quad\bigg{|}\int_{0}^{t}\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}(\theta_{2}-\theta^{(n)}_{2})\rangle\,\mathrm{d}s\bigg{|}\lesssim\int_{0}^{t}\|\theta_{1}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}\|u_{\theta_{1}}\|_{L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})}\|\nabla\Lambda^{-1}(\theta_{2}-\theta^{(n)}_{2})\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}\,\mathrm{d}s$
$\lesssim\int_{0}^{t}\|\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}\|\theta_{1}\|_{H^{-1/2}(\mathbb{T}^{2})}^{1/r}\|\theta_{2}-\theta^{(n)}_{2}\|_{H^{-1/2}(\mathbb{T}^{2})}^{1/r}\left\|\theta_{1}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{1-1/r}\left\|\theta_{2}-\theta^{(n)}_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{1-1/r}\,\mathrm{d}s$
$\lesssim\sup_{t\in[0,T]}\left\|\theta_{1}(t)\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{1/r}\sup_{t\in[0,T]}\left\|\theta_{2}(t)-\theta^{(n)}_{2}(t)\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{1/r}\|\theta_{1}\|_{L^{r}\left([0,T];L^{p}(\mathbb{T}^{2})\right)}$
$\quad\cdot\left(\int_{0}^{t}\left\|\theta_{1}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)^{\frac{1-1/r}{2}}\left(\int_{0}^{t}\left\|\theta_{2}-\theta^{(n)}_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)^{\frac{1-1/r}{2}}\rightarrow 0.$
This completes the proof of (A.3).
Since (A.4) implies that $\theta_{1}$ is sufficiently regular to be taken as a test function, for any $t\in[0,T],$
$\frac{1}{2}\|\theta_{1}(t)\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\left\|\theta_{1}(s)\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\|\theta_{1}(0)\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}+\int_{0}^{t}\langle\Lambda^{\alpha-1/2}\theta_{1},g\rangle\,\mathrm{d}s.$ (A.5)
According to (A.2), (A.3), (A.5) and (A.1), we obtain
$\|\theta_{1}(t)-\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}+\int_{0}^{t}\left\|\theta_{1}-\theta_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$ (A.6)
$\leqslant$ $-\int_{0}^{t}\left\|\theta_{1}-\theta_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s+2\int_{0}^{t}\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}\theta_{1}\rangle+\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}\theta_{2}\rangle\,\mathrm{d}s.$
Using the fact that
$\int_{\mathbb{T}^{2}}\theta(u_{\theta_{2}}-u_{\theta_{1}})\cdot\nabla\Lambda^{-1}(\theta_{2}-\theta_{1})\,\mathrm{d}x=0,$
a rearrangement of the terms yields
$\int_{0}^{t}\langle\theta_{2}u_{\theta_{2}},\nabla\Lambda^{-1}\theta_{1}\rangle+\langle\theta_{1}u_{\theta_{1}},\nabla\Lambda^{-1}\theta_{2}\rangle\,\mathrm{d}s=\int_{0}^{t}\langle(\theta_{2}-\theta_{1})u_{\theta_{1}},\nabla\Lambda^{-1}(\theta_{2}-\theta_{1})\rangle\,\mathrm{d}s.$ (A.7)
Substituting (A.7) into (A.6), and using the estimate
$\quad\left|\int_{0}^{t}\langle(\theta_{2}-\theta_{1})u_{\theta_{1}}\cdot\nabla\Lambda^{-1}(\theta_{2}-\theta_{1})\rangle\,\mathrm{d}s\right|$ (A.8)
$\leqslant C\int_{0}^{t}\|\theta_{2}-\theta_{1}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}\|u_{\theta_{1}}\|_{L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})}\|\nabla\Lambda^{-1}(\theta_{2}-\theta_{1})\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}\,\mathrm{d}s$
$\leqslant C\int_{0}^{t}\|\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}\|\theta_{2}-\theta_{1}\|_{H^{1/p}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$
$\leqslant C\int_{0}^{t}\|\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}\|\theta_{2}-\theta_{1}\|_{H^{-1/2}(\mathbb{T}^{2})}^{2/r}\left\|\theta_{2}-\theta_{1}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2(1-1/r)}\,\mathrm{d}s$
$\leqslant\frac{1}{2}\int_{0}^{t}\left\|\theta_{2}-\theta_{1}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s+C\int_{0}^{t}\|\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}^{r}\left\|\theta_{2}-\theta_{1}\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s,$
we conclude that
$\|\theta_{1}(t)-\theta_{2}(t)\|_{{H^{-1/2}(\mathbb{T}^{2})}}^{2}+\int_{0}^{t}\left\|\theta_{1}-\theta_{2}\right\|_{H^{\alpha-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\lesssim\int_{0}^{t}\|\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}^{r}\left\|\theta_{2}-\theta_{1}\right\|_{H^{-1/2}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s.$
It follows from Grönwall’s inequality that $\theta_{1}-\theta_{2}=0.$
In the next example, we will take $\mathcal{C}^{1}:=\mathbb{X}_{\alpha,\alpha/2+1/4}\cap L^{4}\left([0,T];L^{4}(\mathbb{T}^{2})\right)$ and show that $\mathcal{C}^{1}$ is an $\mathcal{I}$-closure of $\mathcal{C}_{0}^{1}.$
Example A.2.
Let $\theta\in\mathcal{C}^{1}$ with $\mathcal{I}(\theta)<\infty$. Then there exists a sequence $\{\theta^{(n)}\}_{n\geqslant 1}\subset\mathcal{C}_{0}^{1}$ such that $\theta^{(n)}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\alpha/2+1/4}$ and $\mathcal{I}\left(\theta^{(n)}\right)\rightarrow\mathcal{I}(\theta)$ as $n\rightarrow\infty$.
Proof.
For every $n\geqslant 1,$ let $\theta^{(n)}:=\theta\ast\eta_{1/n}$. It suffices to prove
(i) $\theta^{(n)}\in\mathcal{C}_{0}^{1}$,
(ii) $\theta^{(n)}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\alpha/2+1/4}$ as $n\rightarrow\infty,$
(iii) $\mathcal{I}\left(\theta^{(n)}\right)\rightarrow\mathcal{I}(\theta)$.
From the property of convolution kernel $\eta_{\epsilon}$, we deduce that (i) and (ii) hold. Now we proceed to prove (iii). By the lower semi-continuity of $\mathcal{I}$, it suffices to show that $\limsup_{n\rightarrow\infty}\mathcal{I}\left(\theta^{(n)}\right)\leqslant\mathcal{I}(\theta)$.
The fact that $\theta^{(n)}(0)\rightarrow\theta(0)$ in $H^{-1/2}(\mathbb{T}^{2})$ implies $\mathcal{I}_{0}\left(\theta^{(n)}(0)\right)\rightarrow\mathcal{I}_{0}(\theta(0))$. For the dynamic cost, thanks to Proposition 4.2, there exists an element $g\in L^{2}([0,T];H)$ such that
$\partial_{t}\theta=-\Lambda^{2\alpha}\theta-(u_{\theta}\cdot\nabla\theta)+\Lambda^{\alpha+1/2}g$
holds in the sense of (4.4). Taking the convolution with $\eta_{1/n}$, it follows that
$\partial_{t}\theta^{(n)}=-\Lambda^{2\alpha}\theta^{(n)}-(u_{\theta^{(n)}}\cdot\nabla\theta^{(n)})+\Lambda^{\alpha+1/2}g^{(n)}+\Lambda^{\alpha+1/2}R^{(n)},$
where
$R^{(n)}:=\Lambda^{-\alpha-1/2}(u_{\theta^{(n)}}\cdot\nabla\theta^{(n)}-(u_{\theta}\cdot\nabla\theta)^{(n)})=\Lambda^{-\alpha-1/2}\nabla\cdot\left(u_{\theta^{(n)}}\theta^{(n)}-(u_{\theta}\theta)^{(n)}\right).$
We will show that $\left\|R^{(n)}\right\|_{L^{2}([0,T];H)}\rightarrow 0$ as $n\rightarrow\infty$.
Using Cauchy-Schwartz inequality and Calderón-Zygmund inequality, we have
$\|u_{\theta}\theta\|_{L^{2}\left([0,T];L^{2}(\mathbb{T}^{2};\mathbb{R}^{2})\right)}$ $\lesssim\|u_{\theta}\|_{L^{4}\left([0,T];L^{4}(\mathbb{T}^{2};\mathbb{R}^{2})\right)}\|\theta\|_{L^{4}\left([0,T];L^{4}(\mathbb{T}^{2})\right)}$
$\lesssim\|\theta\|_{L^{4}\left([0,T];L^{4}(\mathbb{T}^{2})\right)}^{2}.$
Since $\alpha>1/2$ and $(u_{\theta}\theta)^{(n)}\rightarrow u_{\theta}\theta$ in $L^{2}\left([0,T];L^{2}(\mathbb{T}^{2};\mathbb{R}^{2})\right)$, we find that
$\left\|\Lambda^{-\alpha-1/2}\nabla\cdot\left(u_{\theta}\theta-(u_{\theta}\theta)^{(n)}\right)\right\|_{L^{2}\left([0,T];H\right)}$ $\lesssim\left\|u_{\theta}\theta-(u_{\theta}\theta)^{(n)}\right\|_{L^{2}\left([0,T];H^{1/2-\alpha}(\mathbb{T}^{2};\mathbb{R}^{2})\right)}$
$\lesssim\left\|u_{\theta}\theta-(u_{\theta}\theta)^{(n)}\right\|_{L^{2}\left([0,T];L^{2}(\mathbb{T}^{2};\mathbb{R}^{2})\right)}\rightarrow 0.$
Hence
$R^{1,n}:=\Lambda^{-\alpha-1/2}\nabla\cdot\left(u_{\theta}\theta-(u_{\theta}\theta)^{(n)}\right)\rightarrow 0\quad\text{in}\quad L^{2}\left([0,T];H\right).$
Similarly, we see that $u_{\theta^{(n)}}\theta^{(n)}-u_{\theta}\theta=(u_{\theta^{(n)}}-u_{\theta})\theta^{(n)}+u_{\theta}(\theta^{(n)}-\theta)\rightarrow 0$ in $L^{2}\left([0,T];L^{2}(\mathbb{T}^{2};\mathbb{R}^{2})\right)$. Thus
$R^{2,n}:=\Lambda^{-\alpha-1/2}\nabla\cdot\left(u_{\theta^{(n)}}\theta^{(n)}-u_{\theta}\theta\right)\rightarrow 0\quad\text{in}\quad L^{2}\left([0,T];H\right).$
Therefore, $\left\|R^{(n)}\right\|_{L^{2}([0,T];H)}=\left\|R^{1,n}+R^{2,n}\right\|_{L^{2}([0,T];H)}\rightarrow 0$ as $n\rightarrow\infty$. Since $\theta^{(n)}$ is a weak solution of (1.1) with control $g^{(n)}+R^{(n)}$, we conclude that
$\mathcal{I}_{dyna}\left(\theta^{(n)}\right)\leqslant\left\|g^{(n)}+R^{(n)}\right\|_{L^{2}\left([0,T];H\right)}^{2}\rightarrow\|g\|_{L^{2}\left([0,T];H\right)}^{2}=\mathcal{I}_{dyna}(\theta).$
Remark A.3.
The fact that $\frac{1}{4}+\frac{\alpha}{4}\geqslant\alpha-\frac{1}{2}$ implies that the space $\mathcal{C}^{1}$ is larger than $\mathcal{C}_{0}^{1}$ when taking $r=4$ in the definition of $\mathcal{C}_{0}^{1}$.
Example A.4.
Suppose that $20<\alpha\leqslant 1/$ 2 and $\beta=\alpha/2$. Let
$\mathcal{C}_{0}^{2}=\{\theta\in\mathbb{X}_{\alpha,\alpha/2}:\nabla\theta\in L^{r}\left([0,T];L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})\right),\,\frac{1}{p}+\frac{\alpha}{r}=\alpha,\,p>\frac{1}{\alpha},\,r>0\}.$
For every $g\in L^{2}([0,T];H)$, assume that $\theta_{1}\in\mathcal{C}_{0}^{2}$ and $\theta_{2}\in\mathbb{X}_{\alpha,\alpha/2}$ solve the skeleton equation (1.1) in the sense of Definition 4.1 with the same initial data $\theta_{1}(0)=\theta_{2}(0)\in H$, and suppose further that $\theta_{2}$ satisfy the kinetic energy inequality: for every $t\in[0,T]$,
$\frac{1}{2}\|\theta_{2}(t)\|_{H}^{2}+\int_{0}^{t}\|\theta_{2}(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\frac{1}{2}\left\|\theta_{2}(0)\right\|_{H}^{2}+\int_{0}^{t}\langle\Lambda^{\alpha}\theta_{2},g\rangle\,\mathrm{d}s.$ (A.9)
Then $\theta_{1}=\theta_{2}$ in $\mathbb{X}_{\alpha,\alpha/2}$.
Proof.
This is a direct consequence of [DC06, Theorem 1.1]. We sketch the proof for completeness. For every $t\in[0,T]$, we expand $\|\theta_{1}(t)-\theta_{2}(t)\|_{H}^{2}$ as
$\|\theta_{1}(t)-\theta_{2}(t)\|_{H}^{2}=2\langle\theta_{1}(t),\theta_{2}(t)\rangle+\|\theta_{1}(t)\|_{H}^{2}+\|\theta_{2}(t)\|_{H}^{2}$ (A.10)
and estimate these three terms respectively. For the first term of the right-hand side of (A.10), we need the following auxiliary identity (see [DC06, (2.1)]):
$\quad\int_{0}^{t}\left(2\langle\Lambda^{\alpha}\theta_{1},\Lambda^{\alpha}\theta_{2}\rangle+\langle u_{\theta_{2}}\cdot\nabla\theta_{2},\theta_{1}\rangle+\langle u_{\theta_{1}}\cdot\nabla\theta_{1},\theta_{2}\rangle\right)\,\mathrm{d}s$ (A.11)
$=-\langle\theta_{2}(t),\theta_{1}(t)\rangle+\langle\theta_{2}(0),\theta_{1}(0)\rangle+\int_{0}^{t}\langle\Lambda^{\alpha}(\theta_{2}+\theta_{1}),g\rangle\,\mathrm{d}s.$
Since $\theta_{1}\in\mathcal{C}_{0}^{2}$ is sufficiently regular to be a test function ([Cha06b]), we can apply the $H$-inner product to $\theta_{1}$ to obtain that
$\frac{1}{2}\|\theta_{1}(t)\|_{H}^{2}+\int_{0}^{t}\|\theta_{1}(s)\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s=\frac{1}{2}\left\|\theta_{1}(0)\right\|_{H}^{2}+\int_{0}^{t}\langle\Lambda^{\alpha}\theta_{1},g\rangle\,\mathrm{d}s.$ (A.12)
According to (A.10), (A.11), (A.12) and (A.9), we obtain
$\|\theta_{1}(t)-\theta_{2}(t)\|_{H}^{2}$ $+\int_{0}^{t}\left\|\theta_{1}(s)-\theta_{2}(s)\right\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$ (A.13)
$\leqslant-\int_{0}^{t}\left\|\theta_{1}(s)-\theta_{2}(s)\right\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s+2\int_{0}^{t}\langle u_{\theta_{2}}\cdot\nabla\theta_{2},\theta_{1}\rangle+\langle u_{\theta_{1}}\cdot\nabla\theta_{1},\theta_{2}\rangle\,\mathrm{d}s$
$=-\int_{0}^{t}\left\|\theta_{1}(s)-\theta_{2}(s)\right\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s+2\int_{0}^{t}\langle(u_{\theta_{2}}-u_{\theta_{1}})\cdot\nabla(\theta_{2}-\theta_{1}),\theta_{1}\rangle\,\mathrm{d}s.$
Similar to (A.8), we have the estimate
$\quad\int_{0}^{t}\langle(u_{\theta_{2}}-u_{\theta_{1}})\cdot\nabla(\theta_{2}-\theta_{1}),\theta_{1}\rangle\,\mathrm{d}s$ (A.14)
$\leqslant C\int_{0}^{t}\|u_{\theta_{2}}-u_{\theta_{1}}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2};\mathbb{R}^{2})}\|\nabla\theta_{1}\|_{L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})}\|\theta_{2}-\theta_{1}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}\,\mathrm{d}s$
$\leqslant C\int_{0}^{t}\|\nabla\theta_{1}\|_{L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})}\|\theta_{2}-\theta_{1}\|_{L^{\frac{2p}{p-1}}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$
$\leqslant C\int_{0}^{t}\|\nabla\theta_{1}\|_{L^{p}(\mathbb{T}^{2};\mathbb{R}^{2})}\|\theta_{2}-\theta_{1}\|_{H}^{2/r}\left\|\Lambda^{\alpha}(\theta_{2}-\theta_{1})\right\|_{H}^{2(1-1/r)}\,\mathrm{d}s$
$\leqslant\frac{1}{2}\int_{0}^{t}\left\|\theta_{2}-\theta_{1}\right\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s+C\int_{0}^{t}\|\nabla\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}^{r}\left\|\theta_{2}-\theta_{1}\right\|_{H}^{2}\,\mathrm{d}s,$
In (A.14), we apply Hölder’s inequality in the second line, Calderón-Zygmund inequality in the third line, Gagliardo-Nirenberg interpolation inequality in the fourth line and Young’s equality in the last line. Substituting (A.14) into (A.13), we conclude that
$\|\theta_{1}(t)-\theta_{2}(t)\|_{H}^{2}+\int_{0}^{t}\left\|\theta_{1}(s)-\theta_{2}(s)\right\|_{H^{\alpha}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant C\int_{0}^{t}\|\nabla\theta_{1}\|_{L^{p}(\mathbb{T}^{2})}^{r}\left\|\theta_{2}-\theta_{1}\right\|_{H}^{2}\,\mathrm{d}s.$
It follows from Grönwall’s inequality that $\theta_{1}-\theta_{2}=0.$
In the next example, we will take $\mathcal{C}^{2}:=\mathbb{X}_{\alpha,\alpha/2}\cap L^{4}\left([0,T];H^{1-\alpha,4}(\mathbb{T}^{2})\right)$ and show that $\mathcal{C}^{2}$ is an $\mathcal{I}$-closure of $\mathcal{C}_{0}^{2}.$ Here $H^{1-\alpha,4}(\mathbb{T}^{2})$ is the homogeneous Sobolev space defined by
$H^{s,p}(\mathbb{T}^{2}):=\big{\{}f\in L^{p}(\mathbb{T}^{2}):\text{ there exists some }g\in L^{p}(\mathbb{T}^{2}),f=\Lambda^{-s}g\big{\}}\quad s\geqslant 0,\quad p\geqslant 1,$
with the norm $\|f\|_{H^{s,p}(\mathbb{T}^{2})}:=\|\Lambda^{s}f\|_{L^{p}(\mathbb{T}^{2})}$. We need the following product-type estimate to deal with the nonlinear term.
Lemma A.5.
[Res95, Lemma A.4] Suppose that $s>0$ and $p\in(1,\infty)$. For any $f,g\in C^{\infty}\left(\mathbb{T}^{2}\right)$,
$\left\|\Lambda^{s}(fg)\right\|_{L^{p}(\mathbb{T}^{2})}\leqslant C\left(\|f\|_{L^{p_{1}}(\mathbb{T}^{2})}\left\|\Lambda^{s}g\right\|_{L^{p_{2}}(\mathbb{T}^{2})}+\|g\|_{L^{p_{1}}(\mathbb{T}^{2})}\left\|\Lambda^{s}f\right\|_{L^{p_{2}}(\mathbb{T}^{2})}\right),$
where $p_{1},p_{2}\in(1,\infty)$ satisfy
$\frac{1}{p}=\frac{1}{p_{1}}+\frac{1}{p_{2}}.$
Example A.6.
Let $\theta\in\mathcal{C}^{2}$ with $\mathcal{I}(\theta)<\infty$. Then there exists a sequence $\{\theta^{(n)}\}_{n\geqslant 1}\subset\mathcal{C}_{0}^{2}$ such that $\theta^{(n)}\rightarrow\theta$ in $\mathbb{X}_{\alpha,\alpha/2}$ and $\mathcal{I}\left(\theta^{(n)}\right)\rightarrow\mathcal{I}(\theta)$ as $n\rightarrow\infty$.
Proof.
As in the proof of Example A.2, it is sufficient to show $\left\|R^{(n)}\right\|_{L^{2}([0,T];H)}\rightarrow 0$ as $n\rightarrow\infty$, where
$R^{(n)}:=\Lambda^{-\alpha}(u_{\theta^{(n)}}\cdot\nabla\theta^{(n)}-(u_{\theta}\cdot\nabla\theta)^{(n)})=\Lambda^{-\alpha}\nabla\cdot\left(u_{\theta^{(n)}}\theta^{(n)}-(u_{\theta}\theta)^{(n)}\right).$
Since both Riesz transform and $\Lambda^{1-\alpha}$ are defined by Fourier multipliers, it follows from Calderón-Zygmund inequality that, for all $t\in[0,T]$,
$\|u_{\theta}(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2};\mathbb{R}^{2})}$ $=\|\Lambda^{1-\alpha}R^{\perp}\theta(t)\|_{L^{4}(\mathbb{T}^{2};\mathbb{R}^{2})}=\|R^{\perp}\Lambda^{1-\alpha}\theta(t)\|_{L^{4}(\mathbb{T}^{2};\mathbb{R}^{2})}$
$\leqslant C\|\Lambda^{1-\alpha}\theta(t)\|_{L^{4}(\mathbb{T}^{2})}=C\|\theta(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2})}.$
Using Lemma A.5 with $p_{1}=p_{2}=4$, Cauchy-Schwartz inequality, and Calderón-Zygmund inequality, we have
$\quad\|u_{\theta}\theta\|_{L^{2}\left([0,T];H^{1-\alpha}(\mathbb{T}^{2};\mathbb{R}^{2})\right)}^{2}$
$\lesssim\int_{0}^{T}\|u_{\theta}(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2};\mathbb{R}^{2})}^{2}\|\theta(t)\|_{L^{4}(\mathbb{T}^{2})}^{2}\,\mathrm{d}t+\int_{0}^{T}\|\theta(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2};\mathbb{R}^{2})}^{2}\|u_{\theta}(t)\|_{L^{4}(\mathbb{T}^{2})}^{2}\,\mathrm{d}t$
$\lesssim\bigg{(}\int_{0}^{T}\|u_{\theta}(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2};\mathbb{R}^{2})}^{4}\,\mathrm{d}t\bigg{)}^{1/2}\bigg{(}\int_{0}^{T}\|\theta(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2})}^{4}\,\mathrm{d}t\bigg{)}^{1/2}$
$\quad+\bigg{(}\int_{0}^{T}\|\theta(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2})}^{4}\,\mathrm{d}t\bigg{)}^{1/2}\bigg{(}\int_{0}^{T}\|u_{\theta}(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2};\mathbb{R}^{2})}^{4}\,\mathrm{d}t\bigg{)}^{1/2}$
$\lesssim\int_{0}^{T}\|\theta(t)\|_{H^{1-\alpha,4}(\mathbb{T}^{2})}^{4}\,\mathrm{d}t.$
As $(u_{\theta}\theta)^{(n)}\rightarrow u_{\theta}\theta$ in $L^{2}\left([0,T];H^{1-\alpha}(\mathbb{T}^{2};\mathbb{R}^{2})\right)$, we find that
$R^{1,n}:=\Lambda^{-\alpha}\nabla\cdot\left(u_{\theta}\theta-(u_{\theta}\theta)^{(n)}\right)\rightarrow 0\quad\text{in}\quad L^{2}\left([0,T];H\right).$
Applying Lemma A.5 again, we see that $u_{\theta^{(n)}}\theta^{(n)}-u_{\theta}\theta=(u_{\theta^{(n)}}-u_{\theta})\theta^{(n)}+u_{\theta}(\theta^{(n)}-\theta)\rightarrow 0$ in $L^{2}\left([0,T];H^{1-\alpha}(\mathbb{T}^{2};\mathbb{R}^{2})\right)$. Thus
$R^{2,n}:=\Lambda^{-\alpha}\nabla\cdot\left(u_{\theta^{(n)}}\theta^{(n)}-u_{\theta}\theta\right)\rightarrow 0\quad\text{in}\quad L^{2}\left([0,T];H\right).$
Therefore, $\left\|R^{(n)}\right\|_{L^{2}([0,T];H)}=\left\|R^{1,n}+R^{2,n}\right\|_{L^{2}([0,T];H)}\rightarrow 0$ as $n\rightarrow\infty$. Since $\theta^{(n)}$ is a weak solution of (1.1) with control $g^{(n)}+R^{(n)}$, we conclude that
$\mathcal{I}_{dyna}\left(\theta^{(n)}\right)\leqslant\left\|g^{(n)}+R^{(n)}\right\|_{L^{2}\left([0,T];H\right)}^{2}\rightarrow\|g\|_{L^{2}\left([0,T];H\right)}^{2}=\mathcal{I}_{dyna}(\theta).$
Remark A.7.
The embedding of $H^{1,\frac{1}{1/4+\alpha/2}}(\mathbb{T}^{2})$ into $H^{1-\alpha,4}(\mathbb{T}^{2})$ and the fact that $\frac{1}{4}+\frac{\alpha}{2}+\frac{\alpha}{4}\geqslant\alpha$ imply that the space $\mathcal{C}^{2}$ is larger than $\mathcal{C}_{0}^{2}$ when taking $r=4$ in the definition of $\mathcal{C}_{0}^{2}$.
Appendix B B Large Deviations for Initial Data
In this section, we will verify the large deviations for the initial data of (1.3) under Assumption A. Let $\mu_{\varepsilon,\delta}^{0}:=\mathcal{G}\left(0,\varepsilon Q_{\delta}/2\right)$ be the Gaussian measure on $H^{\alpha-2\beta}(\mathbb{T}^{2})$. Recall that $\mathcal{I}_{0}:H^{\alpha-2\beta}(\mathbb{T}^{2})\rightarrow[0,\infty)$ is defined by (4.2) and $\Lambda_{0}:C_{b}(H^{\alpha-2\beta}_{w}(\mathbb{T}^{2}))\times H^{\alpha-2\beta}(\mathbb{T}^{2})\rightarrow\mathbb{R}$ is defined by (4.5) and (4.6).
Proposition B.1.
Assume that the scaling regime (2.7) holds for $\varepsilon,\delta>0,$ then $\{\mu_{\varepsilon,\delta}^{0}\}_{\varepsilon,\delta>0}$ satisfy the large deviations on $H^{\alpha-2\beta}_{w}(\mathbb{T}^{2})$ with speed $\varepsilon^{-1}$ and rate function $\mathcal{I}_{0}$.
Proof.
Upper bound: Let $X_{1}^{\varepsilon,\delta}$ be an $H^{\alpha-2\beta}(\mathbb{T}^{2})$-valued random variable with the law $\mu_{\varepsilon,\delta}^{0}$. Then for any $\eta\in(0,1/2)$, it follows from (2.7) that
$\varepsilon\log\mathbb{E}\Big{(}\exp\Big{\{}\frac{\eta}{\varepsilon}\left\|X_{1}^{\varepsilon,\delta}\right\|^{2}_{H^{\alpha-2\beta}(\mathbb{T}^{2})}\Big{\}}\Big{)}$ $=\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}\varepsilon\log\mathbb{E}\Big{(}\exp\Big{\{}\frac{\eta}{\varepsilon}|k|^{2\alpha-4\beta}\langle e_{k},X_{1}^{\varepsilon,\delta}\rangle^{2}\Big{\}}\Big{)}$ (B.1)
$=\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}\varepsilon\log(1-\lambda_{\delta,k}^{2}\eta)^{-1/2}$
$\leqslant\varepsilon\sum_{k\in\mathbb{Z}^{2}\backslash\{0\}}\eta\lambda_{\delta,k}^{2}\leqslant C.$
Thanks to (B.1), it suffices to prove the weak large deviations for $\{\mu_{\varepsilon,\delta}^{0}\}$. Let $F_{0}$ be an arbitrary weakly compact set in $H^{\alpha-2\beta}(\mathbb{T}^{2})$. Using the definitions of $\Lambda_{0}$ and $\mu_{\varepsilon,\delta}^{0}$, a straightforward calculation shows that, for any $\phi\in F_{0}$,
$\int_{H^{\alpha-2\beta}(\mathbb{T}^{2})}\exp\left\{\varepsilon^{-1}\Lambda_{0}\left(\psi,\phi\right)\right\}\mu_{\varepsilon,\delta}^{0}(\mathrm{d}\phi)\leqslant 1.$
It follows that
$\mu_{\varepsilon,\delta}^{0}(F_{0})$ $=\int_{H^{\alpha-2\beta}(\mathbb{T}^{2})}I_{\{\phi\in F_{0}\}}\mu_{\varepsilon,\delta}^{0}(\mathrm{d}\phi)\leqslant\sup_{\phi\in F_{0}}\exp\left\{-\varepsilon^{-1}\Lambda_{0}\left(\psi,\phi\right)\right\}=\exp\left\{-\varepsilon^{-1}\inf_{\phi\in F_{0}}\Lambda_{0}\left(\psi,\phi\right)\right\}.$
Consequently, we have
$\varepsilon\log\mu_{\varepsilon,\delta}^{0}(F_{0})\leqslant-\inf_{\phi\in F_{0}}\Lambda_{0}\left(\psi,\phi\right).$
Since $F_{0}$ is weakly compact in $H^{\alpha-2\beta}(\mathbb{T}^{2})$, we can apply [KL99, Appendix 2, Lemma 3.2] to swap the infimum and supremum to get
$\varepsilon\log\mu_{\varepsilon,m}^{0}(F_{0})\leqslant-\sup_{\psi\in H^{\infty}(\mathbb{T}^{2})}\inf_{\phi\in F_{0}}\Lambda_{0}\left(\psi,\phi\right)=-\inf_{\phi\in F_{0}}\sup_{\psi\in H^{\infty}(\mathbb{T}^{2})}\Lambda_{0}\left(\psi,\phi\right)=-\inf_{\phi\in F_{0}}\left\|\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}.$
Lower Bound: We will take $E=H^{\alpha-2\beta}_{w}(\mathbb{T}^{2})$ and $I=\mathcal{I}_{0}$ in Lemma 2.5 to obtain the lower bound of large deviations. For any $\phi\in H^{\alpha-2\beta}(\mathbb{T}^{2})$, set $\pi_{\varepsilon,\delta}^{0,\phi}:=\mathcal{G}\left(\sqrt{Q_{\delta}}\phi,\varepsilon Q_{\delta}/2\right)$ and $\rho^{0}_{\varepsilon,\delta}:=\frac{\mathrm{d}\pi_{\varepsilon,\delta}^{0,\phi}}{\mathrm{d}\mu_{\varepsilon,\delta}^{0}}$. With the help of Cameron-Martin theorem [DPZ14, Theorem 2.23], the relative entropy can be computed by
$\operatorname{Ent}\left(\pi_{\varepsilon,\delta}^{0,\phi}|\mu_{\varepsilon,\delta}^{0}\right)=\int_{H^{\alpha-2\beta}(\mathbb{T}^{2})}\log\rho^{0}_{\varepsilon,\delta}\,\mathrm{d}\pi_{\varepsilon,\delta}^{0,\phi}=\frac{1}{\varepsilon}\left\|Q_{\delta}^{-1/2}\sqrt{Q_{\delta}}\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}=\frac{1}{\varepsilon}\left\|\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}.$
This leads to (2.5). To verify the weak convergence condition (i) in Lemma 2.5, let $X_{2}^{\varepsilon,\delta}$ be an $H^{\alpha-2\beta}(\mathbb{T}^{2})$-valued random variable with the law $\pi_{\varepsilon,\delta}^{0,\phi}$. Then (2.7) implies
$\mathbb{E}\left\|X_{2}^{\varepsilon,\delta}-\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ $=\left\|\sqrt{Q_{\delta}}\phi-\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\mathbb{E}\left\|X_{2}^{\varepsilon,\delta}-\sqrt{Q_{\delta}}\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$
$=\left\|\sqrt{Q}_{\delta}\phi-\phi\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\varepsilon}{2}\operatorname{Tr}\left(Q_{\delta}\right)\rightarrow 0\quad\text{ as }\varepsilon\rightarrow 0.$
Hence $\pi_{\varepsilon,\delta}^{0,\phi}\rightarrow\delta_{\phi}$ weakly in $H_{w}^{\alpha-2\beta}(\mathbb{T}^{2})$. ∎
Acknowledgements
Lin Wang acknowledges the support by National Key R&D Program of China (No.2020YFA0712700), National Natural Science Foundation of China (No. 12090010, 12090014, 12471138) and Key Laboratory of Random Complex Structures and Data Science, Academy of Mathematics and Systems Science, Chinese Academy of Sciences (No. 2008DP173182). Zhengyan Wu acknowledges the support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via IRTG 2235 - Project Number 282638148.